### v1

# Joint Score-Threshold Optimization for Interpretable Risk Assessment Under Partial Supervision

Abstract
Risk assessment tools in healthcare commonly employ point-based scoring systems that map patients to ordinal risk categories via thresholds. While electronic health record (EHR) data presents opportunities for data-driven optimization of these tools, two fundamental challenges impede standard supervised learning: (1) partial supervision arising from intervention-censored outcomes, where only extreme categories can be reliably labeled, and (2) asymmetric misclassification costs that increase with ordinal distance. We propose a mixed-integer programming (MIP) framework that jointly optimizes scoring weights and category thresholds under these constraints. Our approach handles partial supervision through per-instance feasible label sets, incorporates asymmetric distance-aware objectives, and prevents middle-category collapse via minimum threshold gaps. We further develop a CSO relaxation using softplus losses that preserves the ordinal structure while enabling efficient optimization. The framework supports governance constraints including sign restrictions, sparsity, and minimal modifications to incumbent tools, ensuring practical deployability in clinical workflows.
11footnotetext: Center for Health Systems and Policy Modeling, Department of Health Policy and Management, Johns Hopkins University, Baltimore, MD, USA. Email: fganjkh1@jhu.edu22footnotetext: Malone Center for Engineering in Healthcare, Johns Hopkins University, Baltimore, MD, USA33footnotetext: Center for Systems Science and Engineering, Department of Civil and Systems Engineering, Johns Hopkins University, Baltimore, MD, USA. Emails: espring6@jh.edu, kimia@jhu.edu44footnotetext: Department of Physical Medicine and Rehabilitation, School of Medicine, Johns Hopkins University, Baltimore, MD, USA. Email: ehoyer1@jhmi.edu55footnotetext: Johns Hopkins Hospital, Baltimore, MD, USA66footnotetext: Department of Physical Therapy, University of Nevada, Las Vegas, Las Vegas, NV, USA. Email: daniel.young@unlv.edu
1 Introduction
Across healthcare settings, risk assessment is routinely implemented via itemized, checklist-style instruments in which clinicians mark the presence, absence, or severity of predefined risk factors. Each factor contributes a predetermined point value to a total score, and pre-specified thresholds on this score map patients to ordinal risk categories (e.g., Low, Medium, High) that govern downstream clinical workflows such as monitoring frequency or preventive interventions [19, 16, 11, 3, 1]. This linear scoring combined with threshold-based categorization paradigm is widely adopted due to its transparency, auditability, and ease of operationalization [23, 25, 10]. The Johns Hopkins Fall Risk Assessment Tool (JHFRAT) exemplifies this approach, evaluating eight risk factors including age, fall history, medications, mobility, and cognition [19]. Each factor is assigned integer points (e.g., age $\geq$ 80 years: 3 points; impaired mobility: 2 points), and the total score maps to risk categories via thresholds: Low (0-5 points), Medium (6-13 points), and High ($\geq$ 14 points). These categories trigger standardized intervention bundles ranging from universal precautions to intensive monitoring and physical therapy consultations [17, 6].
The increasing availability of granular electronic health record (EHR) data presents an opportunity to optimize such tools through data-driven methods: re-estimating point weights, adjusting thresholds, and encoding deployment constraints while preserving the interpretable structure that clinicians trust [9, 20]. However, the practical realities of clinical deployment introduce two fundamental challenges that violate the assumptions of standard supervised learning approaches. The first challenge is partial supervision with selective labels. In deployment, patient outcomes and clinical interventions are inherently coupled. For fall risk assessment, we can construct high-confidence labels for extremes: patients who experience falls despite interventions (Safe-High) and patients who remain fall-free under minimal intervention (Safe-Low). However, patients receiving intermediate interventions have censored counterfactual outcomes–we cannot observe what would have occurred without intervention. This creates systematic label uncertainty for middle categories, necessitating a partial-label learning framework where each instance $i$ has a feasible set $\mathcal{S}_{i}$ rather than a single label. The second challenge is asymmetric, distance-aware misclassification costs. Operationally, under-triage (classifying a truly high-risk patient as low-risk) can result in preventable adverse events, while over-triage (classifying a low-risk patient as high-risk) merely increases resource utilization. Moreover, the severity of misclassification increases with ordinal distance: misclassifying a high-risk patient as low-risk is more consequential than misclassifying them as medium-risk. Training objectives must therefore incorporate both directional asymmetry and distance-awareness aligned with clinical priorities [7].
Definition 1 (Ordinal Risk Assessment Problem).
Given training data $\{(x_{i},\mathcal{S}_{i},w_{i})\}_{i=1}^{n}$ where $x_{i}\in\mathbb{R}^{p}$ are patient features, $\mathcal{S}_{i}\subseteq\mathcal{K}=\{1,\ldots,K\}$ are feasible label sets indicating possible risk categories, and $w_{i}>0$ are instance weights, the goal is to learn parameters $(\beta,\tau)$ where $\beta\in\mathbb{R}^{p}$ defines a linear scoring function $s(x)=\beta^{\top}x$ and $\tau_{1}<\tau_{2}<\cdots<\tau_{K-1}$ are monotone thresholds, to minimize expected asymmetric ordinal loss subject to interpretability and operational constraints.
While ordinal regression methods have been extensively studied, existing approaches fail to address the unique challenges of clinical risk assessment. Classical ordinal regression methods such as the proportional odds model [14] assume parallel decision boundaries and complete supervision across all categories. Support vector approaches for ordinal regression [4] and threshold models [24] similarly require fully labeled training data. Recent extensions to deep learning for ordinal problems [15, 2] and specialized ranking losses [22] maintain these restrictive assumptions, treating all categories as equally important and all misclassification costs as symmetric. The partial label learning literature addresses scenarios where each instance is associated with a set of candidate labels [5, 27], with disambiguation methods that iteratively refine label assignments [28, 26] and loss-based methods that modify training objectives [8, 13]. However, our setting differs fundamentally in three ways: (1) feasible sets arise from systematic censoring due to clinical interventions rather than random labeling ambiguity, (2) we require preservation of ordinal structure throughout the learning process, and (3) misclassification costs are inherently asymmetric and distance-aware in healthcare applications [12]. Furthermore, while extensions like partial proportional odds models [18] relax the proportional odds assumption, they still assume complete supervision and do not address the selective labeling problem where entire categories may lack reliable labels.
We introduce a constrained mixed-integer optimization framework that addresses these challenges while preserving the interpretable point-score structure essential for clinical adoption. Our approach formalizes the partial-label learning problem for ordinal risk assessment, where each instance $i$ has a feasible label set $\mathcal{S}_{i}$ encoding label uncertainty, and the optimization respects these constraints without fabricating labels for ambiguous cases. We jointly optimize $(\beta,\tau)$ to capture their interdependence under the ordinal structure. We embed misclassification costs that increase with ordinal distance and allow directional asymmetry, directly aligning the training objective with deployment priorities. When middle categories lack labeled examples, we prevent threshold collapse through minimum gap constraints derived from incumbent tools or cross-validation, maintaining clinically meaningful risk stratification. Finally, we derive a constrained score optimization (CSO) relaxation using softplus losses that preserves the asymmetric, distance-aware training signal while enabling efficient optimization for large-scale problems, extending ideas from smooth hinge losses [21] to the ordinal setting with partial supervision.
The remainder of this paper is organized as follows. Section 2 presents our mixed-integer programming formulation and convex relaxation. Section 3 describes experimental evaluation on clinical risk assessment tasks. Section 4 discusses computational considerations, limitations, and deployment aspects. Section 5 concludes with implications for precision medicine.
2 Methods
2.1 Problem Formulation and Notation
We consider the problem of learning a risk stratification tool that maps patient features $x\in\mathbb{R}^{p}$ to ordinal risk categories $\mathcal{K}=\{1,2,\ldots,K\}$ (e.g., Low, Medium, High for $K=3$). The tool consists of a linear scoring function $s(x)=\beta^{\top}x$ and monotone thresholds $\tau_{1}<\tau_{2}<\cdots<\tau_{K-1}$ that partition the score space into $K$ categories. A patient with score $s(x)$ is assigned to category $k$ if $\tau_{k-1}<s(x)\leq\tau_{k}$, where we define $\tau_{0}=$-$\infty$ and $\tau_{K}=+\infty$ for notational convenience.
The key innovation in our formulation is handling partial supervision through feasible label sets. For each training instance $i$, rather than observing a single true label $y_{i}\in\mathcal{K}$, we have a feasible set $\mathcal{S}_{i}\subseteq\mathcal{K}$ that contains the possible true categories. This formulation naturally captures the selective labeling problem in clinical settings: we can confidently identify extreme cases ($\mathcal{S}_{i}=\{1\}$ for definitely low-risk, $\mathcal{S}_{i}=\{K\}$ for definitely high-risk) while middle-risk patients who received preventive interventions have ambiguous labels ($\mathcal{S}_{i}=\{1,2\}$ or $\mathcal{S}_{i}=\{2,3\}$ or even $\mathcal{S}_{i}=\{1,2,3\}$).
2.2 Mixed-Integer Programming Formulation
We formulate the joint learning of scoring weights $\beta\in\mathbb{R}^{p}$ and thresholds $\tau=(\tau_{1},\ldots,\tau_{K-1})$ as a mixed-integer program that handles partial supervision and asymmetric costs. The MIP approach provides exact solutions and naturally incorporates discrete constraints essential for clinical deployment.
Table 1 summarizes the decision variables and their interpretations. The key insight is to use binary indicators $y_{ik}$ to track whether instance $i$’s score exceeds each threshold $\tau_{k}$, and assignment variables $z_{ik}$ to determine the final category.
Table 1: Decision variables and parameters in the MIP formulation
Symbol Domain Description
$\beta$ $\mathbb{R}^{p}$ Scoring weights for features
$\tau_{k}$ $\mathbb{R}$ Threshold between categories $k$ and $k+1$
$s_{i}$ $\mathbb{R}$ Score for instance $i$: $s_{i}=\beta^{\top}x_{i}$
$y_{ik}$ $\{0,1\}$ Indicator: score $s_{i}$ exceeds threshold $\tau_{k}$
$z_{ik}$ $\{0,1\}$ Assignment of instance $i$ to category $k$
$M$ $\mathbb{R}_{+}$ Big-M constant for constraint activation
$\varepsilon$ $\mathbb{R}_{+}$ Margin parameter for numerical stability
$\delta$ $\mathbb{R}_{+}$ Minimum gap between consecutive thresholds
The complete MIP formulation is:
$\min_{\beta,\tau,y,z}\quad$ $\sum_{i=1}^{n}w_{i}\sum_{k=1}^{K}c(k,\mathcal{S}_{i})\cdot z_{ik}$ (1a)
subject to $\sum_{k=1}^{K}z_{ik}=1,\quad\forall i$ (1b)
$\tau_{k}\leq\tau_{k+1}-\delta,\quad\forall k\in\{1,\ldots,K-2\}$ (1c)
$s_{i}-\tau_{k}\geq\varepsilon-M(1-y_{ik}),\quad\forall i,k$ (1d)
$s_{i}-\tau_{k}\leq-\varepsilon+My_{ik},\quad\forall i,k$ (1e)
$z_{i1}\leq 1-y_{i1},\quad\forall i$ (1f)
$z_{ik}\leq y_{i,k-1}-y_{ik},\quad\forall i,k\in\{2,\ldots,K-1\}$ (1g)
$z_{iK}\leq y_{i,K-1},\quad\forall i$ (1h)
$\beta\in\Omega,\quad y_{ik},z_{ik}\in\{0,1\}$ (1i)
Equation (1b) ensures each instance is assigned to exactly one category. Constraint (1c) maintains threshold ordering with minimum separation $\delta$ to prevent degeneracy. Constraints (1d)-(1e) implement the big-M formulation linking scores to threshold crossings: $y_{ik}=1$ if and only if $s_{i}\geq\tau_{k}+\varepsilon$. The margin $\varepsilon>0$ ensures numerical stability and prevents instances from lying exactly on boundaries. Constraints (1f)-(1h) encode the logical relationship between threshold crossings and category assignments: instance $i$ is in category $k$ if it exceeds exactly the first k1k-1 thresholds.
The cost function $c(k,\mathcal{S}_{i})$ in objective (1a) encodes both the distance to the feasible set and directional asymmetry:
$c(k,\mathcal{S}_{i})=\min_{k^{*}\in\mathcal{S}_{i}}\ell(k,k^{*})$ (2)
This formulation assigns to each instance the cost of predicting category $k$ when the true category is the closest feasible one. For singleton feasible sets ($|\mathcal{S}_{i}|=1$), this reduces to standard supervised learning. For larger feasible sets, the optimization naturally selects the most consistent label within the feasible set.
We use an asymmetric ordinal loss that captures clinical priorities:
$\ell(k,k^{*})=\begin{cases}\alpha_{\text{under}}\cdot|k-k^{*}|^{q}&\text{if }k<k^{*}\text{ (under-triage)}\\ \alpha_{\text{over}}\cdot|k-k^{*}|^{q}&\text{if }k>k^{*}\text{ (over-triage)}\\ 0&\text{if }k=k^{*}\end{cases},$ (3)
with $\alpha_{\text{under}}>\alpha_{\text{over}}$ to penalize under-triage more heavily. The exponent $q\geq 1$ controls the growth rate: $q=1$ yields linear penalty growth with ordinal distance, while $q=2$ imposes quadratic penalties for severe misclassifications. In fall risk assessment, we typically set $\alpha_{\text{under}}/\alpha_{\text{over}}\approx 3$ to reflect that missing a high-risk patient is approximately three times worse than over-treating a low-risk patient.
The big-M constraints in (1d)-(1e) require careful selection of $M$ to ensure correctness while maintaining numerical stability. We set $M=2\cdot\max_{i}\|x_{i}\|\cdot B_{\beta}+B_{\tau}$ where $B_{\beta}$ is an upper bound on $\|\beta\|$ and $B_{\tau}$ bounds the threshold range. These bounds can be derived from domain knowledge (e.g., reasonable score ranges in clinical tools) or data-driven estimates. The MIP has $O(nK)$ binary variables and $O(nK)$ constraints, making it tractable for moderate-sized clinical datasets (thousands of patients) using modern solvers. For larger problems, we employ the CSO warm-start strategy described in Section 2.6.
2.3 Non-Degenerate Categories
When middle categories lack labeled examples (common when $\mathcal{S}_{i}\in\{\{1\},\{K\}\}$ for most instances), the optimization may collapse thresholds to eliminate these categories. While this might optimize the training objective, it defeats the clinical purpose of having gradated risk levels for differential intervention. To prevent this while avoiding label fabrication, we impose minimum gap constraints.
For the three-category case (Low/Medium/High), we enforce:
$\tau_{\text{High}}-\tau_{\text{Low}}\geq\Delta_{\min},$ (4)
where $\Delta_{\min}>0$ ensures the medium category maintains meaningful width. This parameter can be set through several approaches:
  1. 1.
    Incumbent-based: Use the gap from existing validated tools (e.g., JHFRAT has $\Delta_{\min}=8$ points between Low and High thresholds)
  2. 2.
    Cross-validation: On a held-out set with known middle-category examples, select $\Delta_{\min}$ that maximizes classification performance
  3. 3.
    Clinical reasoning: Set based on meaningful score differences (e.g., requiring at least 2-3 risk factors to differ between Low and High)
For $K>3$ categories, we can impose similar constraints on consecutive gaps or the total range, depending on which categories lack supervision.
2.4 Governance and Interpretability Constraints
The feasible set $\Omega$ in Model (1i) encodes interpretability and deployment requirements essential for clinical adoption, in particular:
Sign constraints encode clinical knowledge about risk factor directionality. For factors known to increase risk (e.g., age $\geq$ 80, fall history), we enforce $\beta_{j}\geq 0$. For protective factors (e.g., independent mobility), we enforce $\beta_{j}\leq 0$. This prevents counterintuitive weights that would undermine clinical trust:
$\beta_{j}\geq 0\quad\forall j\in\mathcal{J}_{\text{risk}},\quad\beta_{j}\leq 0\quad\forall j\in\mathcal{J}_{\text{protective}}.$ (5)
Sparsity constraints limit cognitive load by bounding the number of active features. Clinical tools typically use 5-10 items for practical usability:
$\sum_{j=1}^{p}u_{j}\leq s_{\max},\quad-Mu_{j}\leq\beta_{j}\leq Mu_{j},\quad u_{j}\in\{0,1\},$ (6)
where binary variables $u_{j}$ indicate whether feature $j$ is active. The constraint ensures that at most $s_{\max}$ features have non-zero weights.
Minimal modification constraints facilitate adoption by limiting changes from incumbent tools. Clinicians are more likely to accept tools that refine rather than replace existing practice:
$|\beta_{j}-\beta_{j}^{(0)}|\leq\Delta_{j}\quad\text{or penalize}\quad\rho\sum_{j=1}^{p}|\beta_{j}-\beta_{j}^{(0)}|,$ (7)
where $\beta^{(0)}$ represents the incumbent weights. The first form enforces hard limits on individual changes, while the second adds a soft penalty to the objective.
Grouping constraints ensure related features are treated consistently. For example, different mobility assessments might be grouped:
$\beta_{j}=\beta_{j^{\prime}}\quad\forall(j,j^{\prime})\in\mathcal{G},$ (8)
where $\mathcal{G}$ contains pairs of features that should have identical weights.
Performance constraints ensure the optimized model meets minimum clinical safety requirements. Using the assignment variables $z_{ik}$ from the MIP formulation, we can directly bound error rates. For instances with known true labels (singleton feasible sets $\mathcal{S}_{i}=\{k_{i}^{*}\}$), we define:
False positive rate constraint (e.g., limiting low-risk patients classified as high-risk):
$\frac{\sum_{i:k_{i}^{*}=1}z_{iK}}{|\{i:k_{i}^{*}=1\}|}\leq\alpha_{\text{FP}}.$ (9)
False negative rate constraint (e.g., limiting high-risk patients classified as low-risk):
$\frac{\sum_{i:k_{i}^{*}=K}z_{i1}}{|\{i:k_{i}^{*}=K\}|}\leq\alpha_{\text{FN}}.$ (10)
More generally, for any true category $k^{*}$ and predicted category $k$, we can constrain:
$\sum_{i\in\mathcal{I}_{k^{*}}}z_{ik}\leq\alpha_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|\quad\text{or}\quad\geq\gamma_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|,$ (11)
where $\mathcal{I}_{k^{*}}=\{i:\mathcal{S}_{i}=\{k^{*}\}\}$ is the set of instances with true label $k^{*}$, and $\alpha_{k^{*},k}$ (or $\gamma_{k^{*},k}$) specifies the maximum (or minimum) acceptable rate.
These constraints are linear in the assignment variables $z_{ik}$ and integrate directly into the MIP formulation (19). They are particularly valuable for ensuring that optimization does not sacrifice critical safety metrics (e.g., sensitivity for high-risk patients) in favor of overall accuracy. Note that these constraints only apply to instances with known ground truth (singleton feasible sets), which aligns with the partial supervision framework.
2.5 Constrained Score Optimization (CSO) Relaxation
The MIP formulation provides exact solutions but faces computational challenges for large datasets. We derive a convex relaxation that preserves the essential structure—ordinal relationships, partial supervision, and asymmetric costs—while enabling efficient optimization via gradient-based methods.
The MIP model uses binary variables to encode category assignments and threshold crossings. The CSO relaxation replaces these discrete decisions with continuous margin-based losses. An important insight is that the ordinal loss decomposes into a sum of boundary crossing penalties, which we can approximate smoothly. Consider predicting category $k$ for an instance with true category $k^{*}$. The ordinal loss $|k-k^{*}|$ equals the number of thresholds between $k$ and $k^{*}$. We can rewrite this as:
  • If $k<k^{*}$: penalty for failing to cross thresholds $\tau_{k},\tau_{k+1},\ldots,\tau_{k^{*}-1}$.
  • If $k>k^{*}$: penalty for incorrectly crossing thresholds $\tau_{k^{*}},\tau_{k^{*}+1},\ldots,\tau_{k-1}$.
This decomposition motivates a margin-based formulation where we penalize violations of desired threshold crossings.
To create a smooth approximation, we replace discrete boundary indicators with smooth softplus penalties. For each instance-threshold pair $(i,k)$, we define
$\phi_{i,k}^{-}$ $=\frac{1}{\alpha}\log\left(1+\exp(\alpha(\tau_{k}-s_{i}))\right)$ (12)
[…]
Similarly, $\phi_{i,k}^{+}=\frac{1}{\alpha}\log(1+\exp(\alpha(s_{i}-\tau_{k})))$ penalizes exceeding threshold $k$ when we shouldn’t. The temperature parameter $\alpha$ controls the approximation quality: as $\alpha\to\infty$, softplus approaches the hinge loss; as $\alpha\to 0$, it becomes linear.
To derive the CSO objective for instance $i$ with feasible set $\mathcal{S}_{i}$, we define the CSO loss function as:
[…]
The weights $\lambda_{j}^{+},\lambda_{j}^{-}>0$ encode the asymmetric importance of each boundary. To reproduce the asymmetric ordinal loss (3), we set:
$\lambda_{j}^{-}$ $=\alpha_{\text{under}}\cdot w_{j}\quad\text{(weight for failing to exceed threshold }j\text{)},$ (14)
$\lambda_{j}^{+}$ $=\alpha_{\text{over}}\cdot w_{j}\quad\text{(weight for incorrectly exceeding threshold }j\text{)}.$ (15)
where $w_{j}$ encodes the positional importance (e.g., $w_{j}=1$ for linear growth, $w_{j}=j$ for quadratic).
The minimization over $k\in\mathcal{S}_{i}$ implements partial supervision: the loss encourages the instance toward the most compatible category within its feasible set. This is convex since it’s the minimum of convex functions over a finite set.
The full CSO formulation can be written as follows.
$\min_{\beta,\tau}\quad$ $\sum_{i=1}^{n}w_{i}\mathcal{L}_{i}(\beta,\tau)+\mu R(\beta)$ (16a)
subject to $\tau_{1}\leq\tau_{2}\leq\cdots\leq\tau_{K-1}$ (16b)
$\tau_{k+1}-\tau_{k}\geq\delta,\quad\forall k\in\{1,\ldots,K-2\}$ (16c)
$\tau_{\text{High}}-\tau_{\text{Low}}\geq\Delta_{\min}\quad\text{(for $K=3$)}$ (16d)
$\beta\in\Omega$ (16e)
where $R(\beta)$ is an optional regularizer (e.g., $\|\beta\|_{2}^{2}$ for ridge, $\|\beta\|_{1}$ for lasso via proximal methods), and $\Omega$ encodes the same governance constraints as the MIP.
The CSO objective is differentiable with respect to both $\beta$ and $\tau$, with gradients
$\frac{\partial\mathcal{L}_{i}}{\partial\beta}$ $=\sum_{k\in\mathcal{S}_{i}^{*}}p_{ik}\left[\sum_{j=1}^{k-1}\lambda_{j}^{+}\sigma(\alpha(s_{i}-\tau_{j}))-\sum_{j=k}^{K-1}\lambda_{j}^{-}\sigma(\alpha(\tau_{j}-s_{i}))\right]x_{i},$ (17)
$\frac{\partial\mathcal{L}_{i}}{\partial\tau_{j}}$ $=\sum_{k\in\mathcal{S}_{i}^{*}}p_{ik}\begin{cases}-\lambda_{j}^{+}\sigma(\alpha(s_{i}-\tau_{j}))&\text{if }j<k\\ \lambda_{j}^{-}\sigma(\alpha(\tau_{j}-s_{i}))&\text{if }j\geq k\end{cases},$ (18)
where $\sigma(t)=1/(1+\exp(-t))$ is the sigmoid function, and $\mathcal{S}_{i}^{*}\subseteq\mathcal{S}_{i}$ contains the categories achieving the minimum in (13). When $|\mathcal{S}_{i}^{*}|>1$, we use subgradients or smooth approximations. The projection operators maintain feasibility:
  • $\Pi_{\Omega}(\beta)$: projects onto governance constraints (e.g., sign, sparsity)
  • $\Pi_{\mathcal{T}}(\tau)$: projects onto ordered thresholds with minimum gaps via isotonic regression with spacing constraints
Proposition 1 (Convexity).
The CSO objective (16) is convex in $(\beta,\tau)$ jointly.
Proof sketch.
The softplus functions $\phi_{i,k}^{\pm}$ are convex. The weighted sum in (13) preserves convexity. The minimum over $k\in\mathcal{S}_{i}$ is convex as the pointwise minimum of convex functions. The constraints form a convex set. ∎
Proposition 2 (Approximation Quality).
As $\alpha\to\infty$, the CSO solution approaches the MIP solution on separable data.
Proposition 2 follows from the softplus converging to the hinge loss, which exactly captures threshold crossing violations.
Algorithm 1 CSO-Guided Mixed-Integer Optimization for Risk Scoring
1:Training data $\{(x_{i},\mathcal{S}_{i},w_{i})\}_{i=1}^{n}$, parameters $\delta,\Delta_{\min},\alpha$, convergence tolerance $\epsilon$
2:Scoring weights $\beta^{*}$, thresholds $\tau^{*}$
3:Phase 1: CSO Warm-Start
4:Initialize $\beta^{(0)},\tau^{(0)}$ from incumbent tool or random
5:for $t=1,2,\ldots$ until convergence do
6:  Compute CSO loss gradients via (13)
7:  $\beta^{(t+1)}\leftarrow\Pi_{\Omega}(\beta^{(t)}-\eta_{t}\nabla_{\beta}\mathcal{L})$
8:  $\tau^{(t+1)}\leftarrow\Pi_{\mathcal{T}}(\tau^{(t)}-\eta_{t}\nabla_{\tau}\mathcal{L})$ $\triangleright$ Maintain ordering and gaps
9:  if $\|\beta^{(t+1)}-\beta^{(t)}\|+\|\tau^{(t+1)}-\tau^{(t)}\|<\epsilon$ then
10:   break
11:  end if
12:end for
[…]
14:Phase 2: MIP Refinement
15:Initialize MIP solver with warm-start solution $(\beta_{\text{warm}},\tau_{\text{warm}})$
16:Set initial bounds: $\beta_{j}\in[\beta_{\text{warm},j}-\rho,\beta_{\text{warm},j}+\rho]$
17:Solve MIP (19) with warm-start and bounds
18:Optional: If integer scores required, add $\beta_{j}\in\mathbb{Z}$ constraints
19:return optimal solution $(\beta^{*},\tau^{*})$
2.6 Two-Phase Optimization: CSO Warm-Start with MIP Refinement
Algorithm 1 presents our two-phase optimization approach. The CSO relaxation in Phase 1 provides a high-quality warm-start solution that significantly accelerates the subsequent MIP solve in Phase 2. This hybrid approach combines the computational efficiency of convex optimization with the exactness of mixed-integer programming. The warm-start not only reduces solve time but also helps avoid poor local optima by providing the MIP solver with a good initial feasible region.
While our formulation allows continuous-valued scoring weights $\beta\in\mathbb{R}^{p}$, practical deployment often benefits from integer-valued scores that are easier for clinicians to compute mentally. Following the approach in [25], we can enforce integrality by adding constraints $\beta_{j}\in\mathbb{Z}$ to the MIP formulation. When combined with appropriate scaling of features, this yields point values similar to existing tools (e.g., 0-5 points per item). The trade-off between interpretability and predictive performance can be explicitly controlled through the granularity of allowed integer values. Our CSO warm-start remains valuable even with integer constraints, as rounding the continuous solution provides an excellent initial integer feasible solution.
3 Experiments
The Johns Hopkins Fall Risk Assessment Tool provides an ideal test case for our framework as it exemplifies all the key challenges we address: partial supervision due to intervention-censored outcomes, asymmetric costs where missing a fall is more harmful than over-treatment, and the need for interpretable integer-valued scores that clinicians can compute manually. In mapping JHFRAT to our general formulation, the feature vector $x_{i}\in\mathbb{R}^{17}$ represents the 17 binary risk factors across seven clinical categories (age, fall history, mobility, etc.), where $x_{ij}=1$ indicates the presence of risk factor $j$ for patient $i$. The scoring function $s(x_{i})=\beta^{\top}x_{i}$ computes the total JHFRAT score, where $\beta\in\mathbb{Z}^{17}_{+}$ contains the integer point values for each risk factor.
The partial supervision challenge manifests clearly in this setting: among 54,209 hospital encounters, the vast majority (80.7%) received targeted fall prevention interventions that obscure their true risk. Following our framework, we construct feasible label sets $\mathcal{S}_{i}$ based on observed outcomes and intervention patterns. Patients who fell despite interventions are confidently labeled as high-risk ($\mathcal{S}_{i}=\{3\}$), while those who remained fall-free without targeted interventions are labeled as low-risk ($\mathcal{S}_{i}=\{1\}$). The remaining patients who received interventions but did not fall have uncertain risk—they might be truly high-risk patients successfully protected by interventions, or lower-risk patients who received unnecessary precautions. Rather than forcing labels on these ambiguous cases, our framework excludes them from the primary optimization while using them for post-hoc validation.
We evaluate our framework on optimizing the Johns Hopkins Fall Risk Assessment Tool using electronic health record data from three Johns Hopkins Health System hospitals: Johns Hopkins Hospital, Johns Hopkins Bayview Medical Center, and Howard County Medical Center. Our dataset comprises 54,209 hospital admissions between March 28, 2022 and October 27, 2023, with complete JHFRAT assessments, intervention records, and fall events. The JHFRAT assessment consists of 17 non-zero coefficient risk factors across seven categories: age, bladder/bowel elimination, cognition, fall history, patient care equipment, medications, and mobility. On average, patients are assessed twice daily with JHFRAT, and our dataset has an average of 1.87 JHFRAT records per patient day.
3.1 Dataset Construction and Partial Labels
A total of 498 hospital encounters in the dataset include at least one fall event, constituting 0.92% of encounters and equating to 1.07 falls per 1,000 patient-days. The rarity of fall events, combined with widespread use of preventive interventions, creates the selective labeling challenge central to our approach. The majority of patients in the data receive one or more fall-prevention interventions during hospitalization. Of the 31 possible fall-prevention interventions, 13 were identified by a team of clinicians as “targeted” for being resource-intensive and capable of meaningfully altering fall risk. Such targeted interventions include the use of bed exit alarms, increased patient rounding, and constant monitoring. We assume that these targeted interventions, due to their intensive nature and established efficacy, can meaningfully obscure underlying fall risk. In contrast, we do not consider general interventions, like education to patients and family about fall risks, to be risk-obscuring as they are universally applied and have minimal direct impact on fall occurrence.
Figure 1 illustrates how we partition the 54,209 patient encounters into three cohorts based on fall events and targeted intervention receipt. The key insight is that intervention receipt creates asymmetric information about true risk: while falls indicate high risk regardless of interventions (since interventions failed to prevent them), the absence of falls has different interpretations depending on intervention status. We partition patient encounters into three cohorts with distinct labeling strategies:
  • Fall cohort: All encounters with at least one fall event, regardless of targeted interventions applied. These encounters are labeled Safe High-Risk.
  • Reference cohort: Encounters with no targeted interventions throughout the encounter, and no fall events. These encounters are labeled Safe Low-Risk.
  • Intervention cohort: Encounters with at least one targeted intervention during the encounter, and no fall events. These encounters are considered to have uncertain risk, and are not labeled for the optimization.
Figure 1: Stratification of inpatient admissions based on fall events and targeted interventions. The intervention cohort was excluded from primary optimization analyses but retained for post-optimization validation.
The extreme imbalance in cohort sizes–with 80.7% of patients in the ambiguous intervention cohort–underscores why standard supervised learning approaches fail in this setting. Simply assigning these 43,799 patients to a risk category based on their JHFRAT scores or fall outcomes would introduce massive label noise, as we would be guessing the counterfactual (what would have happened without interventions) for the majority of our training data. Our partial supervision approach instead acknowledges this uncertainty explicitly, training only on the 10,410 patients (19.3%) for whom we have confident labels while reserving the intervention cohort for post-optimization validation to assess how the optimized models generalize to the ambiguous majority.
3.2 MIP Models
The JHFRAT scale is divided, per current practice, into three risk categories: Low-Risk ($k=1$) for scores less than 6, Moderate-Risk ($k=2$) for scores 6-13, and High-Risk ($k=3$) for scores greater than 13. While our general framework jointly optimizes both scoring weights $\beta$ and thresholds $\tau$, the JHFRAT application presents unique deployment constraints that motivate a modified approach. To maintain compatibility with existing clinical workflows, electronic health record systems, and staff training, we fix the thresholds at their current values ($\tau_{1}=6,\tau_{2}=13$) and focus optimization on the scoring weights $\beta$ alone. This decision reflects a common practical scenario where healthcare systems require incremental refinements to existing tools rather than complete redesigns, as changing risk thresholds would necessitate updates to intervention protocols, clinical guidelines, and quality metrics across the entire health system.
The absence of reliable moderate-risk labels in our dataset further supports this constrained optimization approach. Due to the selective labeling problem, we can confidently identify only extreme cases: patients who fell despite interventions (Safe High-Risk) and those who remained fall-free without targeted interventions (Safe Low-Risk). Patients with moderate JHFRAT scores predominantly received interventions (78.3% in our data), making their true risk indeterminate—they might be correctly classified moderate-risk patients, or they might be misclassified high or low-risk patients. Rather than introducing noise by guessing labels for this ambiguous middle group, we apply singleton feasible sets only to the extreme cases: $\mathcal{S}_{i}=\{1\}$ for all $i$ in the Safe Low-Risk cohort and $\mathcal{S}_{i}=\{3\}$ for all $i$ in the Safe High-Risk cohort.
This configuration represents a special but important case of our general framework where partial supervision is extreme—no labeled examples exist for intermediate categories. The fixed thresholds naturally prevent the middle category collapse problem described in Section 2.3, as the moderate-risk band maintains its width of $\tau_{2}-\tau_{1}=7$ points regardless of the optimization. While this approach doesn’t leverage our framework’s full capability for joint score-threshold learning, it demonstrates its flexibility in handling severely limited supervision scenarios common in healthcare, where intervention effects obscure risk for the majority of patients. Future work could explore threshold optimization using external validation data or prospective trials where moderate-risk labels can be more reliably determined.
We apply three types of governance constraints in our MIP formulations for JHFRAT. Firstly, we apply a uniform lower and upper bound on all model coefficients. Since we only include the risk factors from JHFRAT with positive coefficients, it is reasonable to impose the clinical assumption inherent in the lower bound of 0 that all of these factors can only contribute positively to high fall risk. Furthermore, we limit coefficients to the highest class’s lower bound of 13 for model stability. Secondly, we impose integer constraints on the coefficients to maintain consistency with the format of JHFRAT and ease cognitive burden of manual assessment. Finally, we impose monotonicity constraints for hierarchical risk factors to mantain clinical validity. In particular, the risk factors in the categories of age, medications, and patient care equipment are inherently hierarchical, and we let $P$ denote the set of pairs of risk factor indices corresponding to these pairwise orderings (i.e. the model coefficient for one high-fall risk drug is less than or equal to the coefficient for two or more high-fall risk drugs).
The complete JHFRAT-optimizing MIP formulation is:
$\min_{\beta,\tau,y,z}\quad$ $\sum_{i=1}^{n}w_{i}\sum_{k=1}^{3}c(k,\mathcal{S}_{i})\cdot z_{ik}$ (19a)
subject to $\sum_{k=1}^{3}z_{ik}=1,\quad\forall i$ (19b)
(1d) - (1h) (19c)
$0\leq\beta_{j}\leq 13\quad\forall j$ (19d)
$\beta_{i}\leq\beta_{j}\quad\forall(i,j)\in P$ (19e)
$\beta_{j}\in\mathbb{Z}\quad y_{ik},z_{ik}\in\{0,1\}$ (19f)
We consider four variations to the MIP model above:
  1. 1.
    Symmetric: The ordinal loss for under-triage and over-triage is the same ($\alpha_{under}=\alpha_{over}=1$). Constraints and other hyperparameters follow exactly from Model (19).
  2. 2.
    Asymmetric: The ordinal loss for under-triage is three times that of over-triage ($\alpha_{under}=3,\alpha_{over}=1$). Constraints and other hyperparameters follow exactly from Model (19).
  3. 3.
    False-Positive (FP) constrained: The ordinal loss for under-triage is three times that of over-triage, and all other hyperparameters follow exactly from Model (19). Let $\bar{\beta}$ indicate the JHFRAT model coefficients and let $\bar{z}$ indicate the feasible set binary classification variables imposed deterministically by the MIP formulation with $\bar{\beta}$. We impose the following two constraints, in addition to those in Model (19):
    $\sum_{i\in[n]|0\in S_{i}}z_{i1}$ $\leq\sum_{i\in[n]|0\in S_{i}}\bar{z}_{i1}$ (20a)
    $\sum_{i\in[n]|0\in S_{i}}z_{i2}$ $\leq\sum_{i\in[n]|0\in S_{i}}\bar{z}_{i2}$ (20b)
    These additional performance constraints ensure that the number of Safe Low-Risk labeled encounters classified as Moderate-Risk (20a) and High-Risk (20b) cannot exceed the numbers of such classifications with the JHFRAT model.
  4. 4.
    False-Negative (FN) & False-Positive (FP) constrained: The ordinal loss for under-triage is three times that of over-triage. The ordinal loss for under-triage is three times that of over-triage, and all other hyperparameters follow exactly from Model (19). With the same definition of $\bar{z}$ as above, we impose the following four constraints, in addition to those in Model (19):
    $\sum_{i\in[n]|2\in S_{i}}z_{i0}$ $\leq\sum_{i\in[n]|2\in S_{i}}\bar{z}_{i0}$ (21a)
    $\sum_{i\in[n]|2\in S_{i}}z_{i1}$ $\leq\sum_{i\in[n]|2\in S_{i}}\bar{z}_{i1}$ (21b)
    $\sum_{i\in[n]|0\in S_{i}}z_{i1}$ $\leq 0.05\sum_{i\in[n]|0\in S_{i}}1$ (21c)
    $\sum_{i\in[n]|0\in S_{i}}z_{i2}$ $\leq 0.3\sum_{i\in[n]|0\in S_{i}}1$ (21d)
    The number of Safe High-Risk labeled encounters classified as Low-Risk (21a) and Moderate Risk (21b) cannot exceed the numbers of such classifications with the JHFRAT model. The portion of Safe Low-Risk labeled encounters classified as Moderate-Risk (21c) and High-Risk (21d) cannot exceed given thresholds (30% and 5% respectively). These constraints allow for a small increase from the JHFRAT benchmark of 26.5% and 0.9% for these same categorizations, as seen in Table 2.
We note that this fixed-threshold approach, while more restrictive than full joint optimization, offers several practical advantages: (1) it ensures the optimized tool remains immediately deployable without systemic changes to care protocols, (2) it maintains comparability with historical JHFRAT data and quality metrics, and (3) it reduces the optimization complexity, improving solution stability and interpretability. The resulting problem focuses on finding optimal integer weights $\beta\in\mathbb{Z}^{17}_{+}$ that best separate the extreme risk groups within the established scoring scale, effectively treating this as a constrained parameter refinement problem rather than a full model redesign.
3.3 Experimental Setup
We split the encounters from the combined fall and reference cohorts into an 80% (encounters) training and validation set, and a 20% (encounters) test set, stratified by risk label. We retain the intervention cohort separately as an exogenous dataset for concordance analysis post-optimization. We perform 5-fold cross-validation for training the CSO, and run a hyperparameter gridsearch over the values of $\alpha_{under}$, $\alpha_{over}\in\{0.5,1,2,4\}$ and $\lambda_{1}\in\{0.2,0.5,0.8\}$ (such that $\lambda_{1}=\lambda_{1}^{-}=\lambda_{1}^{+}$ and $\lambda_{2}=1-\lambda_{1}=\lambda_{2}^{-}=\lambda_{2}^{+}$). For each set of hyperparameters, we consider the average coefficients across the cross-validation folds. Each MIP model is trained on the entire training set, utilizing the best feasible rounded coefficients from the CSO gridsearch. If none of the rounded CSO coefficients are feasible for a given MIP model, the JHFRAT coefficients are used for the warm-start.The CSO model is solved with cvxpy, and all MIP variations are solved with Gurobi 12.0.3.
3.4 Baselines and Evaluation Metrics
We compare the outcomes of our approach against the original JHFRAT model with expert-derived weights. Furthermore, we compare the CSO performance with those of the MIP models in order to evaluate the added value of the MIP portion of the optimization pipeline. We utilize the following metrics to compare model performance:
  • Classification Accuracy: We define tight accuracy as the portion of encounters correctly classified into low or high-risk categories. We further define loose accuracy to additionally include all encounters classified as moderate-risk as correctly classified.
  • High-Risk Precision and Recall: To account for the imbalance of the dataset, we rely on the high-risk precision and recall in addition to the accuracy metrics.
  • Binary Classification Performance Measures: We utilize the distribution of scores for the optimized models to generate receiver operating characteristic (ROC) and precision-recall curves. We then utilize the area under each curve, AUROC and AUPRC respectively, as model performance metrics.
4 Results
4.1 Optimized Risk Factor Coefficients
The optimal solutions to the four MIP variations are featured in Table 2, in comparison with the current JHFRAT scoring coefficients. The optimal CSO coefficients for symmetric class loss (hyperparameters $\alpha_{under},\alpha_{over}=1,\lambda_{1}^{-},\lambda_{1}^{+},\lambda_{2}^{-},\lambda_{2}^{+}=1$) are also included as a representative example of CSO results.
Table 2: Risk factor coefficients across baseline and optimized models
Risk Factor JHFRAT CSO Symmetric Asymmetric FP-cap FN&FP-cap
Age 60-69 years 1 1.70 1 0 1 0
Age 70-79 years 2 1.92 2 3 1 1
Age $\geq$ 80 years 3 2.22 2 13 4 2
Incontinence 2 7.33 11 9 4 12
Urgency/frequency 2 3.75 3 6 5 3
Altered awareness 1 7.66 12 9 12 10
Impulsive 2 15.01 13 11 12 7
Lack of understanding 4 5.20 13 12 3 12
One present 1 0.77 0 3 2 2
Two present 2 1.38 1 3 3 6
Three or more present 3 1.56 2 13 3 8
Fall within 6 months 5 2.22 0 6 0 0
One HFRD 3 2.60 2 2 1 0
Two or more HFRD 5 4.04 4 3 3 0
Sedation procedure 7 4.04 4 3 3 0
Requires assistance 2 6.84 11 13 5 8
Unsteady gait 2 6.76 13 10 0 7
Visual/auditory impairment 2 6.92 6 2 9 11
Total Coefficient Sum 49 81.92 100 121 71 89
All optimized models have total coefficient sums greater than the current JHFRAT, and thus we compare feature contributions via their percent contributions to the total coefficient sum. A few trends persist across the model variations. Cognition risk factors feature far more prominently in the MIP models than in the current JHFRAT. Conversely, the history of falls risk factor only persists as non-zero in the asymmetric MIP model, albeit with less proportional importance (7.4%) compared to JHFRAT (10.2%).
4.2 Score Distributions and Categorization
Figure 2 reveals how different optimization objectives reshape the risk score distributions compared to the original JHFRAT. The Kolmogorov-Smirnov (KS) statistic quantifies the maximum vertical separation between the cumulative distribution functions of the Safe Low-Risk and Safe High-Risk cohorts, with larger values indicating better discriminative ability. The associated p-values test the null hypothesis that the two risk groups come from the same distribution—highly significant p-values (all p < 0.001) confirm that all models successfully separate the risk groups, though with varying degrees of separation.
Figure 2: Score distributions for all labeled data (test and train data together)
The distribution patterns directly reflect each model’s clinical optimization strategy:
Original JHFRAT–Conservative but Imprecise: The overlapping distributions (KS = 0.620) show why current practice struggles: many true high-risk patients score in the low-to-moderate range (under-triage), while some low-risk patients score high (over-triage). This overlap drives both missed falls and unnecessary interventions. With 44.8% of eventual fall patients classified as low or moderate risk, critical prevention opportunities are missed.
Symmetric MIP–Maximum Separation for Clear Decision-Making: The stark bimodal distribution (KS = 0.699) eliminates the ambiguous middle ground, pushing patients toward clear low-risk or high-risk classifications. This suits healthcare systems that prefer definitive decisions: either minimal monitoring (score < 6) or full prevention protocols (score > 13). The trade-off is extreme scores (up to 60) that may seem implausible to clinicians, potentially undermining trust. Best for: facilities with binary intervention protocols where middle-ground treatments offer little value.
Asymmetric MIP–Aggressive Fall Prevention: The rightward shift of both distributions reflects the 3:1 penalty for under-triage, effectively lowering the bar for high-risk classification. This achieves 86% sensitivity for falls (vs. 45% in original JHFRAT) but classifies 17% of safe low-risk patients as high-risk. The extreme score range (up to 70) results from the optimizer pushing borderline cases strongly rightward to avoid missing falls. Best for: facilities where fall consequences are severe (e.g., surgical or elderly units) and resources allow broader intervention deployment.
FP-cap MIP–Targeted High-Risk Identification: By constraining false positives to current levels, this model (KS = 0.671) maintains clinician workload while better identifying true high-risk patients. The distribution shows selective rightward movement—pushing true fall patients across the high-risk threshold while keeping most low-risk patients in place. Scores remain clinically plausible (maximum 35). This achieves 45% fall sensitivity with 81% positive predictive value, meaning when it flags high-risk, it’s usually correct. Best for: resource-constrained settings where intervention capacity is fixed and false alarms cause alert fatigue.
FP&FN-cap MIP—Balanced Improvement Within Constraints: The bidirectional constraints produce the most JHFRAT-like distribution while still achieving better separation (KS = 0.683 vs. 0.620). Both cohorts shift toward the extremes but within bounds that maintain operational compatibility. This model essentially "tune” JHFRAT rather than replacing it, achieving 64% fall sensitivity while keeping false positive rates manageable. Best for: conservative healthcare systems requiring evidence that new models won’t disrupt operations before accepting larger changes.
CS–Smooth Risk Stratification: The continuous optimization produces graduated separation without the stark bimodality of MIP solutions. The smooth distribution (KS = 0.700) suggests CSO finds a natural risk continuum rather than forcing binary separation. This may better reflect clinical reality where risk truly is continuous, not discrete. With 66% fall sensitivity and moderate false positives, it balances competing objectives. Best for: initial implementations where extreme changes might face resistance, or as a warm-start for subsequent MIP refinement.
The key insight is that no single distribution is universally "best"–each reflects different clinical priorities and operational constraints. The original JHFRAT’s overlapping distributions explain its poor performance, while our optimized models offer a menu of alternatives: maximum separation (Symmetric MIP) for clear decisions, protective bias (Asymmetric MIP) for high-stakes environments, resource-aware precision (FP-cap) for efficiency, or incremental improvement (FP&FN-cap) for conservative adoption. The extreme scores in unconstrained models, while mathematically optimal, highlight the importance of including clinical face-validity constraints in future deployments.
We calculate the score differentials (MIP score - JHFRAT score) for all encounters in the dataset, as seen in Figure 3. All MIP models results in a positive average score differential, with a heavy upper tail. The asymmetric MIP without any performance constraints results in the most positive skew and the FP-cap MIP results in the least positive skew.
Figure 3: Score differentials (MIP score - JHFRAT score) for all patients in dataset
Skewness in the score differentials is also reflected in the categorization of encounters across the low, moderate and high-risk categories shown in Table 3. Without additional performance constraints, both the symmetric and asymmetric models significantly increase the number of encounters categorized as high-risk in the test data set. The symmetric model results in a 1100% increase in high-risk false positives and a 164% increase in true positives while the asymmetric model has a 1878% increase in high-risk false positives and 200% increase in true positives in the test set. The two models with false positive maximums still achieve notable gains in true positives: 56% and 124% increases from JHFRAT for FP-cap and FN&FP-cap, respectively. Both models also decrease the number of safe high-risk patients categorized as moderate-risk, but increase the number categorized as low-risk. All models except for the asymmetric MIP without performance constraints achieve a small increase in the number of true negative safe low-risk patients identified as low-risk.
Table 3: Test data encounter categorization across baseline and optimized models
Model Encounters Model Categorization
Low-Risk Moderate-Risk High-Risk
Original JHFRAT Safe Low-Risk 1,149 (72.6%) 420 (26.5%) 14 (0.9%)
Safe High-Risk 7 (8.1%) 55 (63.2%) 25 (28.7%)
All 17,007 (31.4%) 30,254 (55.8%) 6,946 (12.8%)
CSO Safe Low-Risk 1,156 (73.0%) 374 (20.5%) 53 (3.3%)
Safe High-Risk 8 (9.2%) 22 (25.3%) 57 (65.5%)
All 15,640 (28.9%) 18,124 (33.43%) 20,443 (37.7%)
Symmetric MIP Safe Low-Risk 1,180 (74.5%) 262 (16.6%) 168 (10.6%)
Safe High-Risk 7 (8.1%) 14 (16.1%) 66 (75.9%)
All 16,207 (29.9%) 10,581 (19.5%) 27,419 (50.58%)
Asymmetric MIP Safe Low-Risk 946 (59.8%) 360 (22.7%) 277 (17.5%)
Safe High-Risk 6 (6.9%) 6 (6.9%) 75 (86.2%)
All 11,278 (20.8%) 9,212 (17.0%) 33,717 (62.2%)
FP-cap MIP Safe Low-Risk 1,258 (79.5%) 316 (20.0%) 9 (0.6%)
Safe High-Risk 12 (13.8%) 36 (41.4%) 39 (44.8%)
All 18,142 (33.5%) 22,892 (42.2%) 13,172 (24.30%)
FP&FN-cap MIP Safe Low-Risk 1,237 (78.1%) 294 (18.6%) 52 (3.3%)
Safe High-Risk 11 (12.6%) 20 (23.0%) 56 (64.4%)
All 18,034 (33.3%) 15,562 (28.7%) 20,611 (38.0%)
4.3 Performance Metrics
Table 4 includes the performance metrics calculated for each model. In general, the optimized models achieve moderate gains in tight accuracy at the expense of loose accuracy. Only the FP-cap MIP improves on the tight accuracy without loss of loose accuracy.
Table 4: Performance metrics on test dataset
Model Accuracy AUC High-Risk Precision High-Risk Recall
Tight Loose AUROC AUPRC
Original JHFRAT 0.70 0.99 0.92 0.51 0.64 0.29
CSO 0.73 0.96 0.94 0.65 0.52 0.66
Symmetric MIP 0.75 0.90 0.94 0.65 0.28 0.76
Asymmetric MIP 0.61 0.83 0.92 0.57 0.21 0.86
FP-cap MIP 0.78 0.99 0.92 0.62 0.81 0.45
FN&FP-cap MIP 0.77 0.96 0.90 0.63 0.52 0.64
All optimization models outperform JHFRAT in area under the precision-recall curve for the testing and training datasets. All models perform similarly with high recall in the testing dataset, with the curves only notably separating at recall below 0.6. In contrast, the optimized models achieve visibly higher precision across all recall points for the training data. The small improvements in AUROC achieved by the models in the training set do not materialize in the testing set.
Figure 4: Receiver Operating Characteristic (ROC) and Precision-Recall performance curves across JHFRAT and five optimization models on training dataset (above) and testing dataset (below).
5 Discussion
5.1 Key Findings
The proposed CSO and mixed-integer programming pipeline achives significant performance gains compared to the current JHFRAT model while offering priority-based tuning options in the hyperparameters. In general, the score distributions and sample categorizations respond as expected to the provided hyperparameters. Utilizing a symmetric distance function for mis-classification loss results in nearly symmetrical mis-classification rates for low-risk and high-risk labeled samples, while the use of an asymmetric 3-to-1 high-risk to low-risk distance function results in higher skew of both classes towards high-risk categorization. The notable exception to expectations is the increase in low-risk false negatives among the two performance-constrained models. The false-negative constraint for the FN&FP-cap MIP should mitigate an increase in such classifications, but this cannot be guaranteed in the testing setting.
While the CSO and MIP models themselves do not inherently discourage moderate-risk classification, when applied in a setting with no such moderate risk labels, the resulting score distributions disaggregate from the mean. This movement away from moderate risk classification can come at the expense of increasing further-class mis-classification, as is the case for the FP-cap and FN&FP-cap models. Nevertheless, these models remain promising options for improving true high-risk classification without drastic increases in the total number of patients classified as high-risk.
5.2 Limitations and Future Work
Several model extensions warrant investigation:
  • Non-linear scoring functions: Extending to tree-based or shallow neural architectures while maintaining interpretability
  • Time-varying risk: Incorporating temporal dynamics in sequential assessment settings
  • Multi-site learning: Federated optimization across institutions with heterogeneous populations
There are a few key limitations with the JHFRAT model experiments. Use of average JHFRAT scores over the encounter for model training results in some non-binary risk score values. These binary risk-score values represent averages over the encounter, but may contribute to extreme values in the optimized coefficients (i.e. coefficients larger than the highest class lower bound) in the absence of enforced upper bounds. The exclusion of all intervention cohort encounters also potentially limits the generalizability of the optimized models. Future work could explore including some patients from the intervention cohort as Safe Moderate-Risk, or apply multiple labels, depending on the number and type of fall prevention interventions applied.
5.3 Clinical Deployment Considerations
It is of particular clinical interest to improve identification of patients at high risk for falls while avoiding an overall increase in risk scores. From a resource perspective, it is impractical and wasteful to provide fall prevention interventions to patients who do not need them. Furthermore, from a clinical perspective, restricting mobility is an effective fall prevention method, but often leads to detrimental functional decline. Therefore, placing such restrictions on low-risk patients is not only unnecessary, but is actively harmful. These types of concerns over false positives extend beyond fall risk assessment. Alert fatigue is a persistent problem that is exacerbated by many false positives in assessment of risks of adverse events such as sepsis. The inclusion of performance constraints and the user-tunable distance function for mis-classification loss serve as tools to control the balance between under-classification and over-classification. Specific constraints on the number of false positives and/or false negatives for the most extreme classes offer assuredness.
In general, the specific features of the proposed MIP formulation that are applied to any given setting can be based on the context of the risk factors, adverse event in question, available data, and current clinical practice. For example, enforcing optimized coefficients to be integers may be unnecessary if assessments are completed automatically and clinicians only observe the final score. Allowing negativity in model coefficients allows for the inclusion of risk-preventative factors in the model.
6 Conclusion
In this work we develop a multi-class mixed integer programming ordinal classification model and a smooth, convex constrained score optimization. These models can apply in a variety of risk assessment contexts, and are particularly well suited for settings with uncertain ground truth or predictive performance requirements compared to an existing model. We utilize the proposed MIP and CSO formulations to conduct data-driven score optimization of the Johns Hopkins Fall Risk Assessment Tool. Across different model variations, we are able to increase identification of patients who fell as high-risk, and control for false positives via model parameters and constraints. Future work can explore refining or adapting the proposed models to other risk assessment settings.

### v2

@@
-Joint Score-Threshold Optimization for Interpretable Risk Assessment Under Partial Supervision
+Joint Score-Threshold Optimization for Interpretable Risk Assessment
 Abstract
 Risk assessment tools in healthcare commonly employ point-based scoring systems that map patients to ordinal risk categories via thresholds.
-While electronic health record (EHR) data presents opportunities for data-driven optimization of these tools, two fundamental challenges impede standard supervised learning: (1) partial supervision arising from intervention-censored outcomes, where only extreme categories can be reliably labeled, and (2) asymmetric misclassification costs that increase with ordinal distance.
-We propose a mixed-integer programming (MIP) framework that jointly optimizes scoring weights and category thresholds under these constraints.
-Our approach handles partial supervision through per-instance feasible label sets, incorporates asymmetric distance-aware objectives, and prevents middle-category collapse via minimum threshold gaps.
-We further develop a CSO relaxation using softplus losses that preserves the ordinal structure while enabling efficient optimization.
-The framework supports governance constraints including sign restrictions, sparsity, and minimal modifications to incumbent tools, ensuring practical deployability in clinical workflows.
-11footnotetext: Center for Health Systems and Policy Modeling, Department of Health Policy and Management, Johns Hopkins University, Baltimore, MD, USA.
-Email: fganjkh1@jhu.edu22footnotetext: Malone Center for Engineering in Healthcare, Johns Hopkins University, Baltimore, MD, USA33footnotetext: Center for Systems Science and Engineering, Department of Civil and Systems Engineering, Johns Hopkins University, Baltimore, MD, USA.
-Emails: espring6@jh.edu, kimia@jhu.edu44footnotetext: Department of Physical Medicine and Rehabilitation, School of Medicine, Johns Hopkins University, Baltimore, MD, USA.
-Email: ehoyer1@jhmi.edu55footnotetext: Johns Hopkins Hospital, Baltimore, MD, USA66footnotetext: Department of Physical Therapy, University of Nevada, Las Vegas, Las Vegas, NV, USA.
-Email: daniel.young@unlv.edu
+While electronic health record (EHR) data presents opportunities for data-driven optimization of these tools, two fundamental challenges impede standard supervised learning: (1) labels are often available only for extreme risk categories due to intervention-censored outcomes, and (2) misclassification cost is asymmetric and increases with ordinal distance.
+We propose a mixed-integer programming (MIP) framework that jointly optimizes scoring weights and category thresholds in the face of these challenges.
+Our approach prevents label-scarce category collapse via threshold constraints, and utilizes an asymmetric, distance-aware objective.
+The MIP framework supports governance constraints, including sign restrictions, sparsity, and minimal modifications to incumbent tools, ensuring practical deployability in clinical workflows.
+We further develop a continuous relaxation of the MIP problem to provide warm-start solutions for more efficient MIP optimization.
+We apply the proposed score optimization framework to a case study of inpatient falls risk assessment using the Johns Hopkins Fall Risk Assessment Tool.
+Keywords: Clinical decision support, fall risk assessment, interpretable machine learning, mixed-integer programming, ordinal classification, risk scoring systems.
 1 Introduction
 Across healthcare settings, risk assessment is routinely implemented via itemized, checklist-style instruments in which clinicians mark the presence, absence, or severity of predefined risk factors.
-Each factor contributes a predetermined point value to a total score, and pre-specified thresholds on this score map patients to ordinal risk categories (e.g., Low, Medium, High) that govern downstream clinical workflows such as monitoring frequency or preventive interventions [19, 16, 11, 3, 1].
-This linear scoring combined with threshold-based categorization paradigm is widely adopted due to its transparency, auditability, and ease of operationalization [23, 25, 10].
-The Johns Hopkins Fall Risk Assessment Tool (JHFRAT) exemplifies this approach, evaluating eight risk factors including age, fall history, medications, mobility, and cognition [19].
-Each factor is assigned integer points (e.g., age $\geq$ 80 years: 3 points; impaired mobility: 2 points), and the total score maps to risk categories via thresholds: Low (0-5 points), Medium (6-13 points), and High ($\geq$ 14 points).
-These categories trigger standardized intervention bundles ranging from universal precautions to intensive monitoring and physical therapy consultations [17, 6].
-The increasing availability of granular electronic health record (EHR) data presents an opportunity to optimize such tools through data-driven methods: re-estimating point weights, adjusting thresholds, and encoding deployment constraints while preserving the interpretable structure that clinicians trust [9, 20].
-However, the practical realities of clinical deployment introduce two fundamental challenges that violate the assumptions of standard supervised learning approaches.
-The first challenge is partial supervision with selective labels.
-In deployment, patient outcomes and clinical interventions are inherently coupled.
-For fall risk assessment, we can construct high-confidence labels for extremes: patients who experience falls despite interventions (Safe-High) and patients who remain fall-free under minimal intervention (Safe-Low).
-However, patients receiving intermediate interventions have censored counterfactual outcomes–we cannot observe what would have occurred without intervention.
-This creates systematic label uncertainty for middle categories, necessitating a partial-label learning framework where each instance $i$ has a feasible set $\mathcal{S}_{i}$ rather than a single label.
-The second challenge is asymmetric, distance-aware misclassification costs.
-Operationally, under-triage (classifying a truly high-risk patient as low-risk) can result in preventable adverse events, while over-triage (classifying a low-risk patient as high-risk) merely increases resource utilization.
-Moreover, the severity of misclassification increases with ordinal distance: misclassifying a high-risk patient as low-risk is more consequential than misclassifying them as medium-risk.
-Training objectives must therefore incorporate both directional asymmetry and distance-awareness aligned with clinical priorities [7].
-Definition 1 (Ordinal Risk Assessment Problem).
-Given training data $\{(x_{i},\mathcal{S}_{i},w_{i})\}_{i=1}^{n}$ where $x_{i}\in\mathbb{R}^{p}$ are patient features, $\mathcal{S}_{i}\subseteq\mathcal{K}=\{1,\ldots,K\}$ are feasible label sets indicating possible risk categories, and $w_{i}>0$ are instance weights, the goal is to learn parameters $(\beta,\tau)$ where $\beta\in\mathbb{R}^{p}$ defines a linear scoring function $s(x)=\beta^{\top}x$ and $\tau_{1}<\tau_{2}<\cdots<\tau_{K-1}$ are monotone thresholds, to minimize expected asymmetric ordinal loss subject to interpretability and operational constraints.
-While ordinal regression methods have been extensively studied, existing approaches fail to address the unique challenges of clinical risk assessment.
+Each factor contributes a predetermined integer point value to a total score, and pre-specified integer thresholds map patient scores to ordinal risk categories (e.g., Low, Medium, High) that govern downstream clinical workflows such as monitoring frequency or preventive interventions [20, 18, 12, 5, 1].
+The restriction to integer coefficients and thresholds is not merely conventional but essential for practical deployment, enabling rapid bedside assessment and facilitating clinician trust through cognitive transparency.
+This point-based scoring paradigm is widely adopted due to its transparency and ease of use: clinicians can compute risk scores through simple mental calculations without requiring calculators or computational tools [24, 27, 11].
+However, many widely used point-based risk scoring tools were developed based on clinical reasoning rather than data-driven evidence, and warrant updating for current clinical practice and changing populations [28, 16].
+The increasing availability of granular electronic health record (EHR) data presents an opportunity to optimize such tools through data-driven methods: re-estimating integer point weights, adjusting integer thresholds, and encoding deployment constraints while preserving both the interpretable structure and computational simplicity that clinicians trust [10, 21].
+The practical realities of clinical deployment introduce two fundamental challenges that violate the assumptions of standard supervised learning approaches.
+The first challenge is a frequent lack of ground-truth labels for patients’ level of risk.
+In deployment, patient outcomes and clinical interventions are inherently coupled, preventing clear causal analysis between risk factors and outcomes.
+The second challenge is with promoting model assuredness and alignment with clinical priorities.
+Unconstrained supervised learning can result in unexpected outcomes that do not align with the realities of available resources or clinical objectives.
+Operationally, under-triage (classifying a truly high-risk patient as low-risk) can result in preventable adverse events with serious consequences, while over-triage (classifying a low-risk patient as high-risk) can result in resource inefficiency, alert fatigue, and worse patient outcomes.
+The relative severity of under- versus over-triage depends on the specific adverse event, available interventions, and patient population.
+While some existing models can balance this tradeoff via model hyperparameters, they are not equipped to safely control the resulting risk score distributions.
+We introduce a constrained mixed-integer optimization framework for patient risk classification that addresses these challenges while preserving the interpretable point-score structure essential for clinical adoption.
+Our approach handles settings where labels exist only for extreme risk categories.
+In the extreme case (which applies in our clinical application), intervention-censored outcomes prevent reliable labeling of middle-risk categories entirely.
+We can confidently identify only the extremes: patients who experienced adverse events despite interventions (high-risk) and those who remained event-free without resource-intensive interventions (low-risk).
+Patients who received interventions but remained event-free have maximally uncertain labels, as it is not possible to fully determine whether interventions prevented adverse events or were unnecessarily applied.
+For these ambiguous cases, we do not force a label and exclude them from training.
+We retain these cases for validation, comparing the distributions of scores between risk categories across models.
+We embed misclassification costs that increase with ordinal distance and allow directional asymmetry, directly aligning the training objective with deployment priorities.
+When middle categories lack labeled examples, we prevent threshold collapse through minimum gap constraints derived from incumbent tools or cross-validation, maintaining clinically meaningful risk stratification.
+Finally, we extend ideas from smooth hinge losses [23] to the ordinal setting with our derived constrained score optimization (CSO) relaxation, which uses softplus losses to preserve the asymmetric, distance-aware training signal while enabling efficient optimization for large-scale problems.
+We develop a joint model optimization pipeline with MIP and CSO, and apply this methodology to a case study of improving the Johns Hopkins Fall Risk Assessment Tool (JHFRAT) for assessing inpatient fall risk.
+The remainder of this paper is organized as follows.
+Section 2 highlights related work in the area of ordinal risk assessment with regression methods and constrained optimization.
+Section 3 presents our mixed-integer programming formulation and convex relaxation.
+Section 4 describes experimental evaluation on clinical risk assessment tasks.
+Section 6 discusses computational considerations, limitations, and deployment aspects.
+Section 7 concludes with implications for precision medicine.
+2 Related Work
+Ordinal regression methods have been extensively studied, but existing approaches fail to address the unique challenges of clinical risk assessment.
 Classical ordinal regression methods such as the proportional odds model [15] assume parallel decision boundaries and complete supervision across all categories.
@@
 The partial label learning literature addresses scenarios where each instance is associated with a set of candidate labels [7, 30], with disambiguation methods that iteratively refine label assignments [31, 29] and loss-based methods that modify training objectives [9, 14].
-However, our setting differs fundamentally in three ways: (1) feasible sets arise from systematic censoring due to clinical interventions rather than random labeling ambiguity, (2) we require preservation of ordinal structure throughout the learning process, and (3) misclassification costs are inherently asymmetric and distance-aware in healthcare applications [12].
+However, our setting differs fundamentally in three ways: (1) feasible sets arise from systematic censoring due to clinical interventions rather than random labeling ambiguity, (2) misclassification costs are inherently asymmetric and distance-aware in healthcare applications, and (3) we require preservation of ordinal structure throughout the learning process [13].
 Furthermore, while extensions like partial proportional odds models [19] relax the proportional odds assumption, they still assume complete supervision and do not address the selective labeling problem where entire categories may lack reliable labels.
-We introduce a constrained mixed-integer optimization framework that addresses these challenges while preserving the interpretable point-score structure essential for clinical adoption.
-Our approach formalizes the partial-label learning problem for ordinal risk assessment, where each instance $i$ has a feasible label set $\mathcal{S}_{i}$ encoding label uncertainty, and the optimization respects these constraints without fabricating labels for ambiguous cases.
-We jointly optimize $(\beta,\tau)$ to capture their interdependence under the ordinal structure.
-We embed misclassification costs that increase with ordinal distance and allow directional asymmetry, directly aligning the training objective with deployment priorities.
-When middle categories lack labeled examples, we prevent threshold collapse through minimum gap constraints derived from incumbent tools or cross-validation, maintaining clinically meaningful risk stratification.
-Finally, we derive a constrained score optimization (CSO) relaxation using softplus losses that preserves the asymmetric, distance-aware training signal while enabling efficient optimization for large-scale problems, extending ideas from smooth hinge losses [21] to the ordinal setting with partial supervision.
-The remainder of this paper is organized as follows.
-Section 2 presents our mixed-integer programming formulation and convex relaxation.
-Section 3 describes experimental evaluation on clinical risk assessment tasks.
-Section 4 discusses computational considerations, limitations, and deployment aspects.
-Section 5 concludes with implications for precision medicine.
-2 Methods
-2.1 Problem Formulation and Notation
+We build upon the literature for constrained mixed integer optimization methods as an alternative to ordinal regression.
+Early work employing integer programming methods for classification includes that of Uney and Turkay [26] and Bertsimas and Shioda [2].
+Both works solve supervised multi-class classification problems by separating different classes into polyhedral regions of the data space.
+Bertsimas and Shioda also solve a separate linear regression problem, which employs unsupervised classification to create different linear models across heterogeneous subpopulations.
+Billiet et al.
+[3] formulate a risk-score weighting binary classification linear programming model based on soft-margin support vector machines.
+The Interval Coded Scoring linear program presented in that work utilizes similar model constraints as integer programming classification, but avoids the use of integer or binary variables via the SVM-based objective.
+Dedieu et al.
+[8] address the computational complexity of scaling integer programming classification models by solving a series of smaller relaxed MIP subproblems.
+Ustun and Rudin [27] present the RiskSlim model, which optimizes integer risk model coefficients.
+They present both a nonlinear objective model and a surrogate linear program model.
+Among the current literature, their work is the most similar to our methodology, though their model applies only to binary risk classification.
+Furthermore, we use a discrete objective function to enforce class-based ordinal misclassification.
+3 Proposed Models
 We consider the problem of learning a risk stratification tool that maps patient features $x\in\mathbb{R}^{p}$ to ordinal risk categories $\mathcal{K}=\{1,2,\ldots,K\}$ (e.g., Low, Medium, High for $K=3$).
-The tool consists of a linear scoring function $s(x)=\beta^{\top}x$ and monotone thresholds $\tau_{1}<\tau_{2}<\cdots<\tau_{K-1}$ that partition the score space into $K$ categories.
+Following the clinical scoring system paradigm [27], the tool consists of integer-valued scoring weights $\beta\in\mathbb{Z}^{p}$ and integer thresholds $\tau=(\tau_{1},\ldots,\tau_{K-1})\in\mathbb{Z}^{K-1}$ with $\tau_{1}<\tau_{2}<\cdots<\tau_{K-1}$.
+The integer restriction enables clinicians to compute risk scores $s(x)=\beta^{\top}x$ through mental arithmetic without computational aids, a critical requirement for bedside deployment and cognitive transparency.
 A patient with score $s(x)$ is assigned to category $k$ if $\tau_{k-1}<s(x)\leq\tau_{k}$, where we define $\tau_{0}=$-$\infty$ and $\tau_{K}=+\infty$ for notational convenience.
-The key innovation in our formulation is handling partial supervision through feasible label sets.
-For each training instance $i$, rather than observing a single true label $y_{i}\in\mathcal{K}$, we have a feasible set $\mathcal{S}_{i}\subseteq\mathcal{K}$ that contains the possible true categories.
-This formulation naturally captures the selective labeling problem in clinical settings: we can confidently identify extreme cases ($\mathcal{S}_{i}=\{1\}$ for definitely low-risk, $\mathcal{S}_{i}=\{K\}$ for definitely high-risk) while middle-risk patients who received preventive interventions have ambiguous labels ($\mathcal{S}_{i}=\{1,2\}$ or $\mathcal{S}_{i}=\{2,3\}$ or even $\mathcal{S}_{i}=\{1,2,3\}$).
+Given training data $\{(x_{i},k_{i}^{*},w_{i})\}_{i\in\mathcal{I}}$ where $x_{i}\in\mathbb{R}^{p}$ are patient features, $k_{i}^{*}\in\mathcal{K}$ are known category labels, and $w_{i}>0$ are instance weights, our goal is to learn integer parameters $(\beta,\tau)\in\mathbb{Z}^{p}\times\mathbb{Z}^{K-1}$ that minimize expected asymmetric ordinal loss subject to interpretability and operational constraints.
+Our setting considers the possibility that some patient samples will not have associated risk labels.
+Without imposition of assumptions or causal analysis of the effects of interventions, these samples must be excluded from training to maintain the integrity of the training labels.
+The key challenge then becomes partial supervision at the category level: while each included training instance $i\in\mathcal{I}$ has a single known label $k_{i}^{*}$, certain categories $k\in\mathcal{K}$ may have no labeled training examples.
+Let $\mathcal{I}_{k}=\{i\in\mathcal{I}:k_{i}^{*}=k\}$ denote the set of training instances with known label $k$.
+In the extreme case where only the most extreme categories have labeled examples (i.e., $\mathcal{I}_{k}=\emptyset$ for intermediate categories), the optimization may collapse thresholds to eliminate unlabeled categories.
+Section 3.2 addresses this degeneracy through minimum threshold gap constraints.
 3.1 Mixed-Integer Programming Formulation
-We formulate the joint learning of scoring weights $\beta\in\mathbb{R}^{p}$ and thresholds $\tau=(\tau_{1},\ldots,\tau_{K-1})$ as a mixed-integer program that handles partial supervision and asymmetric costs.
-The MIP approach provides exact solutions and naturally incorporates discrete constraints essential for clinical deployment.
-Table 1 summarizes the decision variables and their interpretations.
-The key insight is to use binary indicators $y_{ik}$ to track whether instance $i$’s score exceeds each threshold $\tau_{k}$, and assignment variables $z_{ik}$ to determine the final category.
+We formulate the joint learning of integer scoring weights $\beta\in\mathbb{Z}^{p}$ and integer thresholds $\tau=(\tau_{1},\ldots,\tau_{K-1})\in\mathbb{Z}^{K-1}$ as a mixed-integer program that handles partial supervision and asymmetric costs.
+The MIP approach provides exact solutions and naturally incorporates both the integrality constraints and other discrete constraints essential for clinical deployment.
+Table 1 summarizes the decision variables and parameters in our formulation.
+The key insight is to use assignment variables $z_{ik}$ to determine the category assignment for each instance, with big-M constraints enforcing that assigned categories respect the threshold boundaries.
+The complete MIP formulation is given by:
+$\min_{\beta,\tau,z}\quad$ $\sum_{i\in\mathcal{I}}w_{i}\sum_{k=1}^{K}\ell(k,k_{i}^{*})\cdot z_{ik}$ (1a)
+subject to $\sum_{k=1}^{K}z_{ik}=1,\quad\forall i\in\mathcal{I}$ (1b)
+$\tau_{k}\leq\tau_{k+1}-\delta,\quad\forall k\in\{1,\ldots,K-2\}$ (1c)
+$s_{i}=\beta^{T}x_{i}\quad\forall i\in\mathcal{I}$ (1d)
+$M(z_{ik}-1)+\tau_{k-1}+\varepsilon\leq s_{i},$
+$\qquad\forall i\in\mathcal{I},\,k\in\{1,\ldots,K\}$ (1e)
+$s_{i}\leq M(1-z_{ik})+\tau_{k},$
+$\qquad\forall i\in\mathcal{I},\,k\in\{1,\ldots,K\}$ (1f)
+$\beta\in\Omega\cap\mathbb{Z}^{p}$ (1g)
+$z\in\Theta\cap\{0,1\}^{K|\mathcal{I}|}$ (1h)
+$\tau\in\mathbb{Z}^{K-1}$ (1i)
 Table 1: Decision variables and parameters in the MIP formulation
 Symbol Domain Description
-$\beta$ $\mathbb{R}^{p}$ Scoring weights for features
-$\tau_{k}$ $\mathbb{R}$ Threshold between categories $k$ and $k+1$
-$s_{i}$ $\mathbb{R}$ Score for instance $i$: $s_{i}=\beta^{\top}x_{i}$
-$y_{ik}$ $\{0,1\}$ Indicator: score $s_{i}$ exceeds threshold $\tau_{k}$
+$\beta$ $\mathbb{Z}^{p}$ Scoring weights
+$\tau_{k}$ $\mathbb{Z}$ Threshold between categories $k$ and $k+1$
+$s_{i}$ $\mathbb{Z}$ Score for instance $i$: $s_{i}=\beta^{\top}x_{i}$
 $z_{ik}$ $\{0,1\}$ Assignment of instance $i$ to category $k$
-$M$ $\mathbb{R}_{+}$ Big-M constant for constraint activation
-$\varepsilon$ $\mathbb{R}_{+}$ Margin parameter for numerical stability
-$\delta$ $\mathbb{R}_{+}$ Minimum gap between consecutive thresholds
-The complete MIP formulation is:
-$\min_{\beta,\tau,y,z}\quad$ $\sum_{i=1}^{n}w_{i}\sum_{k=1}^{K}c(k,\mathcal{S}_{i})\cdot z_{ik}$ (1a)
-subject to $\sum_{k=1}^{K}z_{ik}=1,\quad\forall i$ (1b)
-$\tau_{k}\leq\tau_{k+1}-\delta,\quad\forall k\in\{1,\ldots,K-2\}$ (1c)
-$s_{i}-\tau_{k}\geq\varepsilon-M(1-y_{ik}),\quad\forall i,k$ (1d)
-$s_{i}-\tau_{k}\leq-\varepsilon+My_{ik},\quad\forall i,k$ (1e)
-$z_{i1}\leq 1-y_{i1},\quad\forall i$ (1f)
-$z_{ik}\leq y_{i,k-1}-y_{ik},\quad\forall i,k\in\{2,\ldots,K-1\}$ (1g)
-$z_{iK}\leq y_{i,K-1},\quad\forall i$ (1h)
-$\beta\in\Omega,\quad y_{ik},z_{ik}\in\{0,1\}$ (1i)
+$M$ $\mathbb{R}_{+}$ Big-M constant
+$\varepsilon$ $\mathbb{R}_{+}$ Margin parameter
+$\delta$ $\mathbb{Z}_{+}$ Min. gap between thresholds
 Equation (1b) ensures each instance is assigned to exactly one category.
-Constraint (1c) maintains threshold ordering with minimum separation $\delta$ to prevent degeneracy.
-Constraints (1d)-(1e) implement the big-M formulation linking scores to threshold crossings: $y_{ik}=1$ if and only if $s_{i}\geq\tau_{k}+\varepsilon$.
-The margin $\varepsilon>0$ ensures numerical stability and prevents instances from lying exactly on boundaries.
-Constraints (1f)-(1h) encode the logical relationship between threshold crossings and category assignments: instance $i$ is in category $k$ if it exceeds exactly the first k1k-1 thresholds.
-The cost function $c(k,\mathcal{S}_{i})$ in objective (1a) encodes both the distance to the feasible set and directional asymmetry:
-$c(k,\mathcal{S}_{i})=\min_{k^{*}\in\mathcal{S}_{i}}\ell(k,k^{*})$ (2)
-This formulation assigns to each instance the cost of predicting category $k$ when the true category is the closest feasible one.
-For singleton feasible sets ($|\mathcal{S}_{i}|=1$), this reduces to standard supervised learning.
-For larger feasible sets, the optimization naturally selects the most consistent label within the feasible set.
-We use an asymmetric ordinal loss that captures clinical priorities:
-$\ell(k,k^{*})=\begin{cases}\alpha_{\text{under}}\cdot|k-k^{*}|^{q}&\text{if }k<k^{*}\text{ (under-triage)}\\ \alpha_{\text{over}}\cdot|k-k^{*}|^{q}&\text{if }k>k^{*}\text{ (over-triage)}\\ 0&\text{if }k=k^{*}\end{cases},$ (3)
-with $\alpha_{\text{under}}>\alpha_{\text{over}}$ to penalize under-triage more heavily.
+Constraint (1c) maintains threshold ordering with minimum separation $\delta>0$ to prevent degeneracy, as detailed in Section 3.2.
+Constraints (1e)–(1f) implement the big-M formulation enforcing threshold bounds to assigned categories: if $z_{ik}=1$, then $\tau_{k-1}+\varepsilon\leq s_{i}\leq\tau_{k}$.
+The margin $\varepsilon>0$ ensures numerical stability and prevents ambiguity when scores lie exactly on a boundary.
+The domain constraints in (1g) and (1h) encapsulate governance and performance constraints on $\beta$ and $z$, respectively, that are described in Sections 3.3 and 3.4.
+We use an asymmetric ordinal loss function $\ell(k,k^{*})$ in objective (1a) that captures the differential consequences of misclassification:
+$\ell(k,k^{*})=\begin{cases}\omega_{\text{under}}\cdot|k-k^{*}|^{q}&\text{if }k<k^{*}\\ \omega_{\text{over}}\cdot|k-k^{*}|^{q}&\text{if }k>k^{*}\\ 0&\text{if }k=k^{*}\end{cases},$ (2)
+where $\omega_{\text{under}}>\omega_{\text{over}}$ reflects that under-triage typically carries more severe consequences than over-triage in healthcare settings.
 The exponent $q\geq 1$ controls the growth rate: $q=1$ yields linear penalty growth with ordinal distance, while $q=2$ imposes quadratic penalties for severe misclassifications.
-In fall risk assessment, we typically set $\alpha_{\text{under}}/\alpha_{\text{over}}\approx 3$ to reflect that missing a high-risk patient is approximately three times worse than over-treating a low-risk patient.
-The big-M constraints in (1d)-(1e) require careful selection of $M$ to ensure correctness while maintaining numerical stability.
+The specific ratio $\omega_{\text{under}}/\omega_{\text{over}}$ should be determined based on the clinical context, weighing the relative costs of missed adverse events versus unnecessary interventions.
+The big-M constraints in (1e)–(1f) require careful selection of $M$ to ensure correctness while maintaining numerical stability.
 We set $M=2\cdot\max_{i}\|x_{i}\|\cdot B_{\beta}+B_{\tau}$ where $B_{\beta}$ is an upper bound on $\|\beta\|$ and $B_{\tau}$ bounds the threshold range.
@@
 For larger problems, we employ the CSO warm-start strategy described in Section 3.6.
-2.3 Non-Degenerate Categories
-When middle categories lack labeled examples (common when $\mathcal{S}_{i}\in\{\{1\},\{K\}\}$ for most instances), the optimization may collapse thresholds to eliminate these categories.
-While this might optimize the training objective, it defeats the clinical purpose of having gradated risk levels for differential intervention.
-To prevent this while avoiding label fabrication, we impose minimum gap constraints.
-For the three-category case (Low/Medium/High), we enforce:
-$\tau_{\text{High}}-\tau_{\text{Low}}\geq\Delta_{\min},$ (4)
-where $\Delta_{\min}>0$ ensures the medium category maintains meaningful width. This parameter can be set through several approaches:
-  1. 1.
-    Incumbent-based: Use the gap from existing validated tools (e.g., JHFRAT has $\Delta_{\min}=8$ points between Low and High thresholds)
-  2. 2.
-    Cross-validation: On a held-out set with known middle-category examples, select $\Delta_{\min}$ that maximizes classification performance
-  3. 3.
-    Clinical reasoning: Set based on meaningful score differences (e.g., requiring at least 2-3 risk factors to differ between Low and High)
+3.2 Handling Missing Category Labels
+In the extreme partial supervision setting, where certain categories lack any labeled examples, the optimization faces a fundamental challenge: without training instances to anchor middle categories, the objective may collapse thresholds to eliminate these categories entirely.
+While this maximizes performance on labeled extremes, it defeats the clinical purpose of having gradated risk levels that guide differential interventions.
+Consider the common scenario where only extreme cases can be reliably labeled: $k^{*}_{i}\in\{1,K\}$ for all training instances.
+An optimizer focused solely on separating these extremes might set $\tau_{1}\approx\tau_{2}\approx\cdots\approx\tau_{K-1}$, effectively collapsing all middle categories.
+This would achieve optimal separation on training data but produce a binary classifier rather than the desired ordinal tool.
+Therefore, to maintain clinically meaningful risk stratification without fabricating labels for unlabeled categories, we impose minimum gap constraints per (1c), where $\delta>0$ ensures all categories maintain meaningful width.
+This parameter can be set through several approaches.
+In the case where we are optimizing an existing tool, we can utilize the existing tool’s threshold gaps.
+For example, if the incumbent assessment tool has a gap of 8 points between low-risk and high-risk thresholds, we set $\delta=8$.
+On a held-out set with known middle-category examples, we can select $\delta$ via cross-validation to maximize classification performance.
+Alternatively, we can apply clinical reasoning to set meaningful score differences (e.g., requiring at least 2–3 risk factors to differ between Low and High categories).
 For $K>3$ categories, we can impose similar constraints on consecutive gaps or the total range, depending on which categories lack supervision.
-2.4 Governance and Interpretability Constraints
-The feasible set $\Omega$ in Model (1i) encodes interpretability and deployment requirements essential for clinical adoption, in particular:
-Sign constraints encode clinical knowledge about risk factor directionality.
-For factors known to increase risk (e.g., age $\geq$ 80, fall history), we enforce $\beta_{j}\geq 0$.
-For protective factors (e.g., independent mobility), we enforce $\beta_{j}\leq 0$.
-This prevents counterintuitive weights that would undermine clinical trust:
-$\beta_{j}\geq 0\quad\forall j\in\mathcal{J}_{\text{risk}},\quad\beta_{j}\leq 0\quad\forall j\in\mathcal{J}_{\text{protective}}.$ (5)
-Sparsity constraints limit cognitive load by bounding the number of active features. Clinical tools typically use 5-10 items for practical usability:
+3.3 Governance Constraints
+The feasible set $\Omega$ in (1g) encodes interpretability and deployment requirements essential for clinical adoption.
+These constraints ensure that optimized models remain transparent, clinically valid, and aligned with existing practice.
+Clinical knowledge about risk factor directionality can be encoded through sign restrictions on the weights.
+For factors known to increase risk, we enforce $\beta_{j}\geq 0$, while we enforce $\beta_{j}\leq 0$ for protective factors.
+This prevents counterintuitive weights that would undermine clinical trust, $\beta_{j}\geq 0\quad\forall j\in\mathcal{J}_{\text{risk}},$ and $\beta_{j}\leq 0\quad\forall j\in\mathcal{J}_{\text{protective}}.$
+To limit cognitive load during manual assessment, we can bound the number of active features through sparsity constraints.
+Clinical tools typically use 5–10 items for practical usability.
+Introducing binary variables $u_{j}$ to indicate whether feature $j$ is active, we enforce:
 $\sum_{j=1}^{p}u_{j}\leq s_{\max},\quad-Mu_{j}\leq\beta_{j}\leq Mu_{j},\quad u_{j}\in\{0,1\},$ (3)
-where binary variables $u_{j}$ indicate whether feature $j$ is active. The constraint ensures that at most $s_{\max}$ features have non-zero weights.
-Minimal modification constraints facilitate adoption by limiting changes from incumbent tools. Clinicians are more likely to accept tools that refine rather than replace existing practice:
-$|\beta_{j}-\beta_{j}^{(0)}|\leq\Delta_{j}\quad\text{or penalize}\quad\rho\sum_{j=1}^{p}|\beta_{j}-\beta_{j}^{(0)}|,$ (7)
-where $\beta^{(0)}$ represents the incumbent weights. The first form enforces hard limits on individual changes, while the second adds a soft penalty to the objective.
-Grouping constraints ensure related features are treated consistently. For example, different mobility assessments might be grouped:
-$\beta_{j}=\beta_{j^{\prime}}\quad\forall(j,j^{\prime})\in\mathcal{G},$ (8)
-where $\mathcal{G}$ contains pairs of features that should have identical weights.
-Performance constraints ensure the optimized model meets minimum clinical safety requirements.
-Using the assignment variables $z_{ik}$ from the MIP formulation, we can directly bound error rates.
-For instances with known true labels (singleton feasible sets $\mathcal{S}_{i}=\{k_{i}^{*}\}$), we define:
-False positive rate constraint (e.g., limiting low-risk patients classified as high-risk):
-$\frac{\sum_{i:k_{i}^{*}=1}z_{iK}}{|\{i:k_{i}^{*}=1\}|}\leq\alpha_{\text{FP}}.$ (9)
-False negative rate constraint (e.g., limiting high-risk patients classified as low-risk):
-$\frac{\sum_{i:k_{i}^{*}=K}z_{i1}}{|\{i:k_{i}^{*}=K\}|}\leq\alpha_{\text{FN}}.$ (10)
-More generally, for any true category $k^{*}$ and predicted category $k$, we can constrain:
-$\sum_{i\in\mathcal{I}_{k^{*}}}z_{ik}\leq\alpha_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|\quad\text{or}\quad\geq\gamma_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|,$ (11)
-where $\mathcal{I}_{k^{*}}=\{i:\mathcal{S}_{i}=\{k^{*}\}\}$ is the set of instances with true label $k^{*}$, and $\alpha_{k^{*},k}$ (or $\gamma_{k^{*},k}$) specifies the maximum (or minimum) acceptable rate.
-These constraints are linear in the assignment variables $z_{ik}$ and integrate directly into the MIP formulation (19).
+where $s_{\max}$ bounds the total number of features with non-zero weights.
+Clinical adoption is often facilitated by limiting changes from incumbent tools, as clinicians are more likely to accept refinements rather than wholesale replacements of existing practice.
+Let $\beta^{(0)}$ represent the incumbent weights.
+We can either enforce hard limits on individual changes or add a soft penalty to the objective:
+$|\beta_{j}-\beta_{j}^{(0)}|\leq\Delta_{j}\quad\text{or penalize}\quad\rho\sum_{j=1}^{p}|\beta_{j}-\beta_{j}^{(0)}|.$ (4)
+The first form constrains each weight individually, while the second adds an $L_{1}$ penalty that encourages overall similarity to the incumbent model.
+When features are inherently ordinal, monotonicity constraints can enforce the proper ordinal relationship.
+For example, the risk coefficient for taking two high fall risk medications should be at least as large as the coefficient for taking one high fall risk medication.
+Let $\mathcal{G}$ denote the set of feature pairs that should be ordered: $\beta_{j}\leq\beta_{j^{\prime}}\quad\forall(j,j^{\prime})\in\mathcal{G}.$
+All of these interpretability constraints can be made to be linear and thus integrated directly into the MIP formulation, allowing optimization solvers to efficiently navigate the constrained solution space while maintaining clinical validity.
+The specific constraints deployed depend on the clinical context.
+Sign restrictions are nearly universal, sparsity is important when manual computation is required, minimal modification eases adoption of updated tools, and ordering ensures consistency and robustness against training data noise when features measure related constructs.
+3.4 Performance Constraints
+Performance constraints, encoded via the domain constraints in (1h), ensure the optimized model meets minimum safety and efficacy requirements.
+Using the assignment variables $z_{ik}$ from the MIP formulation, we can directly constrain classification error rates and bound operational metrics like sensitivity and specificity.
+For false positive rate control, we can limit the fraction of low-risk patients misclassified into higher risk categories.
+For example, to ensure that no more than $\xi_{\text{FPR}}$ fraction of truly low-risk patients are classified as high-risk:
+$\frac{\sum_{i:k_{i}^{*}=1}z_{iK}}{|\{i:k_{i}^{*}=1\}|}\leq\xi_{\text{FPR}}.$ (5)
+Similarly, for false negative rate control, we can limit the fraction of high-risk patients misclassified as low-risk. To ensure sensitivity of at least $1-\xi_{\text{FNR}}$ for high-risk detection:
+$\frac{\sum_{i:k_{i}^{*}=K}z_{i1}}{|\{i:k_{i}^{*}=K\}|}\leq\xi_{\text{FNR}}.$ (6)
+More generally, for any true category $k^{*}$ and predicted category $k$, we can impose upper or lower bounds on misclassification rates:
+$\sum_{i\in\mathcal{I}_{k^{*}}}z_{ik}\leq u_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|\quad\text{or}\quad\geq l_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|,$ (7)
+where $\mathcal{I}_{k^{*}}=\{i:\mathcal{S}_{i}=\{k^{*}\}\}$ denotes instances with known true label $k^{*}$, and $u_{k^{*},k}$ (or $l_{k^{*},k}$) specifies the maximum (or minimum) acceptable rate.
+These constraints are linear in the assignment variables and integrate directly into the MIP formulation (1).
 They are particularly valuable for ensuring that optimization does not sacrifice critical safety metrics (e.g., sensitivity for high-risk patients) in favor of overall accuracy.
-Note that these constraints only apply to instances with known ground truth (singleton feasible sets), which aligns with the partial supervision framework.
+In practice, performance constraints can be derived from incumbent tool performance (ensuring the new model does not regress on key metrics) or from external regulatory or quality requirements.
 3.5 Constrained Score Optimization (CSO) Relaxation
 The MIP formulation provides exact solutions but faces computational challenges for large datasets.
-We derive a convex relaxation that preserves the essential structure—ordinal relationships, partial supervision, and asymmetric costs—while enabling efficient optimization via gradient-based methods.
+We derive a convex relaxation that preserves the essential structures of ordinal relationships, partial supervision, and asymmetric costs while enabling efficient optimization via gradient-based methods.
 The MIP model uses binary variables to encode category assignments and threshold crossings.
@@
 An important insight is that the ordinal loss decomposes into a sum of boundary crossing penalties, which we can approximate smoothly.
-Consider predicting category $k$ for an instance with true category $k^{*}$.
-The ordinal loss $|k-k^{*}|$ equals the number of thresholds between $k$ and $k^{*}$.
-We can rewrite this as:
+Consider predicting category $k$ for an instance with true category $k^{*}$. The ordinal loss $|k-k^{*}|$ equals the number of thresholds between $k$ and $k^{*}$. We can rewrite this as:
   • If $k<k^{*}$: penalty for failing to cross thresholds $\tau_{k},\tau_{k+1},\ldots,\tau_{k^{*}-1}$.
@@
 To create a smooth approximation, we replace discrete boundary indicators with smooth softplus penalties. For each instance-threshold pair $(i,k)$, we define
-$\phi_{i,k}^{-}$ $=\frac{1}{\alpha}\log\left(1+\exp(\alpha(\tau_{k}-s_{i}))\right)$ (12)
-[…]
-Similarly, $\phi_{i,k}^{+}=\frac{1}{\alpha}\log(1+\exp(\alpha(s_{i}-\tau_{k})))$ penalizes exceeding threshold $k$ when we shouldn’t.
-The temperature parameter $\alpha$ controls the approximation quality: as $\alpha\to\infty$, softplus approaches the hinge loss; as $\alpha\to 0$, it becomes linear.
-To derive the CSO objective for instance $i$ with feasible set $\mathcal{S}_{i}$, we define the CSO loss function as:
-[…]
-The weights $\lambda_{j}^{+},\lambda_{j}^{-}>0$ encode the asymmetric importance of each boundary. To reproduce the asymmetric ordinal loss (3), we set:
-$\lambda_{j}^{-}$ $=\alpha_{\text{under}}\cdot w_{j}\quad\text{(weight for failing to exceed threshold }j\text{)},$ (14)
-$\lambda_{j}^{+}$ $=\alpha_{\text{over}}\cdot w_{j}\quad\text{(weight for incorrectly exceeding threshold }j\text{)}.$ (15)
-where $w_{j}$ encodes the positional importance (e.g., $w_{j}=1$ for linear growth, $w_{j}=j$ for quadratic).
-The minimization over $k\in\mathcal{S}_{i}$ implements partial supervision: the loss encourages the instance toward the most compatible category within its feasible set.
-This is convex since it’s the minimum of convex functions over a finite set.
-The full CSO formulation can be written as follows.
+$\phi_{i,k}^{-}$ $=\log\left(1+\exp(\alpha^{-}(\tau_{k}-s_{i}))\right)$ (8)
+$\approx\begin{cases}0&\text{if }s_{i}\gg\tau_{k}\\ \alpha^{-}(\tau_{k}-s_{i})&\text{if }s_{i}\ll\tau_{k}\\ \log(1+\exp(\alpha^{-}))&\text{if }s_{i}=\tau_{k}\end{cases}.$
+Similarly, $\phi_{i,k}^{+}=\log(1+\exp(\alpha^{+}(s_{i}-\tau_{k})))$ penalizes exceeding threshold $k$ when we should not.
+The temperature parameters $\alpha^{-}$ and $\alpha^{+}$ control the steepness of the linear penalty for the score falling on the wrong side of the threshold.
+To derive the CSO objective for instance $i$ with category label $k_{i}^{*}$, we define the CSO loss function as:
+$\mathcal{L}_{i}(\beta,\tau)=\min_{k\in\mathcal{S}_{i}}\sum_{j=1}^{k_{i}^{*}-1}\lambda_{j}^{+}\phi_{i,j}^{+}+\sum_{j=k_{i}^{*}}^{K-1}\lambda_{j}^{-}\phi_{i,j}^{-},$ (9)
+where the weights $\lambda_{j}^{+},\lambda_{j}^{-}>0$ encode the asymmetric importance of each boundary. The full CSO formulation can be written as follows:
 $\min_{\beta,\tau}\quad$ $\sum_{i=1}^{n}w_{i}\mathcal{L}_{i}(\beta,\tau)+\mu R(\beta)$ (10a)
@@
 $\tau_{k+1}-\tau_{k}\geq\delta,\quad\forall k\in\{1,\ldots,K-2\}$ (10c)
-$\tau_{\text{High}}-\tau_{\text{Low}}\geq\Delta_{\min}\quad\text{(for $K=3$)}$ (16d)
-$\beta\in\Omega$ (16e)
+$\beta\in\Omega$ (10d)
 where $R(\beta)$ is an optional regularizer (e.g., $\|\beta\|_{2}^{2}$ for ridge, $\|\beta\|_{1}$ for lasso via proximal methods), and $\Omega$ encodes the same governance constraints as the MIP.
-The CSO objective is differentiable with respect to both $\beta$ and $\tau$, with gradients
-$\frac{\partial\mathcal{L}_{i}}{\partial\beta}$ $=\sum_{k\in\mathcal{S}_{i}^{*}}p_{ik}\left[\sum_{j=1}^{k-1}\lambda_{j}^{+}\sigma(\alpha(s_{i}-\tau_{j}))-\sum_{j=k}^{K-1}\lambda_{j}^{-}\sigma(\alpha(\tau_{j}-s_{i}))\right]x_{i},$ (17)
-$\frac{\partial\mathcal{L}_{i}}{\partial\tau_{j}}$ $=\sum_{k\in\mathcal{S}_{i}^{*}}p_{ik}\begin{cases}-\lambda_{j}^{+}\sigma(\alpha(s_{i}-\tau_{j}))&\text{if }j<k\\ \lambda_{j}^{-}\sigma(\alpha(\tau_{j}-s_{i}))&\text{if }j\geq k\end{cases},$ (18)
-where $\sigma(t)=1/(1+\exp(-t))$ is the sigmoid function, and $\mathcal{S}_{i}^{*}\subseteq\mathcal{S}_{i}$ contains the categories achieving the minimum in (13).
-When $|\mathcal{S}_{i}^{*}|>1$, we use subgradients or smooth approximations.
-The projection operators maintain feasibility:
-  • $\Pi_{\Omega}(\beta)$: projects onto governance constraints (e.g., sign, sparsity)
-  • $\Pi_{\mathcal{T}}(\tau)$: projects onto ordered thresholds with minimum gaps via isotonic regression with spacing constraints
-Proposition 1 (Convexity).
-The CSO objective (16) is convex in $(\beta,\tau)$ jointly.
-Proof sketch.
-The softplus functions $\phi_{i,k}^{\pm}$ are convex.
+The CSO objective is smooth and thus differentiable everywhere with respect to both $\beta$ and $\tau$.
+Furthermore, the CSO objective (10) is convex in $(\beta,\tau)$ jointly because the softplus functions $\phi_{i,k}^{\pm}$ are convex.
 The weighted sum in (9) preserves convexity.
-The minimum over $k\in\mathcal{S}_{i}$ is convex as the pointwise minimum of convex functions.
-The constraints form a convex set.
-∎
-Proposition 2 (Approximation Quality).
-As $\alpha\to\infty$, the CSO solution approaches the MIP solution on separable data.
-Proposition 2 follows from the softplus converging to the hinge loss, which exactly captures threshold crossing violations.
-Algorithm 1 CSO-Guided Mixed-Integer Optimization for Risk Scoring
-1:Training data $\{(x_{i},\mathcal{S}_{i},w_{i})\}_{i=1}^{n}$, parameters $\delta,\Delta_{\min},\alpha$, convergence tolerance $\epsilon$
-2:Scoring weights $\beta^{*}$, thresholds $\tau^{*}$
-3:Phase 1: CSO Warm-Start
-4:Initialize $\beta^{(0)},\tau^{(0)}$ from incumbent tool or random
-5:for $t=1,2,\ldots$ until convergence do
-6:  Compute CSO loss gradients via (13)
-7:  $\beta^{(t+1)}\leftarrow\Pi_{\Omega}(\beta^{(t)}-\eta_{t}\nabla_{\beta}\mathcal{L})$
-8:  $\tau^{(t+1)}\leftarrow\Pi_{\mathcal{T}}(\tau^{(t)}-\eta_{t}\nabla_{\tau}\mathcal{L})$ $\triangleright$ Maintain ordering and gaps
-9:  if $\|\beta^{(t+1)}-\beta^{(t)}\|+\|\tau^{(t+1)}-\tau^{(t)}\|<\epsilon$ then
-10:   break
-11:  end if
-12:end for
-[…]
-14:Phase 2: MIP Refinement
-15:Initialize MIP solver with warm-start solution $(\beta_{\text{warm}},\tau_{\text{warm}})$
-16:Set initial bounds: $\beta_{j}\in[\beta_{\text{warm},j}-\rho,\beta_{\text{warm},j}+\rho]$
-17:Solve MIP (19) with warm-start and bounds
-18:Optional: If integer scores required, add $\beta_{j}\in\mathbb{Z}$ constraints
-19:return optimal solution $(\beta^{*},\tau^{*})$
+The differentiability and convexity of the objective, along with the convexity of the feasible set formed by the constraints allows for efficient solution of CSO using commercial nonlinear solvers.
 3.6 Two-Phase Optimization: CSO Warm-Start with MIP Refinement
-Algorithm 1 presents our two-phase optimization approach.
-The CSO relaxation in Phase 1 provides a high-quality warm-start solution that significantly accelerates the subsequent MIP solve in Phase 2.
-This hybrid approach combines the computational efficiency of convex optimization with the exactness of mixed-integer programming.
-The warm-start not only reduces solve time but also helps avoid poor local optima by providing the MIP solver with a good initial feasible region.
-While our formulation allows continuous-valued scoring weights $\beta\in\mathbb{R}^{p}$, practical deployment often benefits from integer-valued scores that are easier for clinicians to compute mentally.
-Following the approach in [25], we can enforce integrality by adding constraints $\beta_{j}\in\mathbb{Z}$ to the MIP formulation.
-When combined with appropriate scaling of features, this yields point values similar to existing tools (e.g., 0-5 points per item).
-The trade-off between interpretability and predictive performance can be explicitly controlled through the granularity of allowed integer values.
-Our CSO warm-start remains valuable even with integer constraints, as rounding the continuous solution provides an excellent initial integer feasible solution.
+The CSO relaxation can provide a high-quality warm-start solution that significantly accelerates the subsequent MIP solve.
+The major challenge with implementing this warm-start approach is ensuring MIP feasibility of the warm-start solution.
+Simple rounding can be applied to adhere to integrality constraints for the risk model coefficients.
+Any continuous linear governance constraints from the MIP can be added to the CSO formulation without compromising differentiability and convexity.
+Performance constraints, on the other hand, cannot be directly added to CSO due to their reliance on the MIP binary variables.
+However, hyperparameter tuning of the CSO $\alpha$ and $\lambda$ values can be used to find solutions which satisfy these constraints.
+This hybrid approach of using CSO and MIP combines the computational efficiency of convex optimization with the exactness of mixed-integer programming.
+The warm-start reduces solve time and helps avoid poor local optima by providing the MIP solver with a good initial feasible region.
 4 Experiments
-The Johns Hopkins Fall Risk Assessment Tool provides an ideal test case for our framework as it exemplifies all the key challenges we address: partial supervision due to intervention-censored outcomes, asymmetric costs where missing a fall is more harmful than over-treatment, and the need for interpretable integer-valued scores that clinicians can compute manually.
-In mapping JHFRAT to our general formulation, the feature vector $x_{i}\in\mathbb{R}^{17}$ represents the 17 binary risk factors across seven clinical categories (age, fall history, mobility, etc.), where $x_{ij}=1$ indicates the presence of risk factor $j$ for patient $i$.
+The Johns Hopkins Fall Risk Assessment Tool provides an ideal test case for our framework as it exemplifies all the key challenges we address: labels available only for extreme categories due to intervention-censored outcomes, asymmetric costs where missing a fall is more harmful than over-treatment, and the need for interpretable integer-valued scores that clinicians can compute manually.
+In mapping JHFRAT to our general formulation, the feature vector $x_{i}\in\{0,1\}^{17}$ represents the 17 binary risk factors across seven clinical categories (age, fall history, mobility, etc.), where $x_{ij}=1$ indicates the presence of risk factor $j$ for patient $i$.
 The scoring function $s(x_{i})=\beta^{\top}x_{i}$ computes the total JHFRAT score, where $\beta\in\mathbb{Z}^{17}_{+}$ contains the integer point values for each risk factor.
-The partial supervision challenge manifests clearly in this setting: among 54,209 hospital encounters, the vast majority (80.7%) received targeted fall prevention interventions that obscure their true risk.
-Following our framework, we construct feasible label sets $\mathcal{S}_{i}$ based on observed outcomes and intervention patterns.
-Patients who fell despite interventions are confidently labeled as high-risk ($\mathcal{S}_{i}=\{3\}$), while those who remained fall-free without targeted interventions are labeled as low-risk ($\mathcal{S}_{i}=\{1\}$).
-The remaining patients who received interventions but did not fall have uncertain risk—they might be truly high-risk patients successfully protected by interventions, or lower-risk patients who received unnecessary precautions.
-Rather than forcing labels on these ambiguous cases, our framework excludes them from the primary optimization while using them for post-hoc validation.
 We evaluate our framework on optimizing the Johns Hopkins Fall Risk Assessment Tool using electronic health record data from three Johns Hopkins Health System hospitals: Johns Hopkins Hospital, Johns Hopkins Bayview Medical Center, and Howard County Medical Center.
@@
 On average, patients are assessed twice daily with JHFRAT, and our dataset has an average of 1.87 JHFRAT records per patient day.
+This study was approved by the Johns Hopkins University Institutional Review Board: IRB00395329, Functional measures in the hospital and patient outcomes.
 4.1 Dataset Construction and Partial Labels
 A total of 498 hospital encounters in the dataset include at least one fall event, constituting 0.92% of encounters and equating to 1.07 falls per 1,000 patient-days.
-The rarity of fall events, combined with widespread use of preventive interventions, creates the selective labeling challenge central to our approach.
 The majority of patients in the data receive one or more fall-prevention interventions during hospitalization.
@@
 We assume that these targeted interventions, due to their intensive nature and established efficacy, can meaningfully obscure underlying fall risk.
-In contrast, we do not consider general interventions, like education to patients and family about fall risks, to be risk-obscuring as they are universally applied and have minimal direct impact on fall occurrence.
-Figure 1 illustrates how we partition the 54,209 patient encounters into three cohorts based on fall events and targeted intervention receipt.
-The key insight is that intervention receipt creates asymmetric information about true risk: while falls indicate high risk regardless of interventions (since interventions failed to prevent them), the absence of falls has different interpretations depending on intervention status.
+In contrast, we do not consider general interventions, like educating patients to wait/call for assistance prior to mobilizing, to be risk-obscuring as they are universally applied and have minimal direct impact on fall occurrence.
+The partial supervision challenge manifests clearly in this setting: among 54,209 hospital encounters, the vast majority (80.7%) received targeted fall prevention interventions that obscure their true risk.
+The rarity of fall events, combined with widespread use of preventive interventions, creates the selective labeling challenge central to our approach.
+For patients who experienced any falls, we consider only the JHFRAT assessments prior to the first fall event.
+To obtain a medically representative sample for the study cohort, we include patients who fell between the 2nd and 21st days of the encounter, or whose overall length of stay was between 2 and 21 full days.
+We exclude all encounters with fewer than two JHFRAT assessments during the included encounter time.
+After applying this filtering, the study cohort includes 46,991 patient encounters.
+Figure 1 illustrates how we partition the 46,991 filtered patient encounters into three cohorts based on fall events and targeted intervention receipt.
+The key insight is that intervention receipt creates asymmetric information about true risk: while falls indicate high risk regardless of interventions (since interventions failed to prevent them), the absence of falls can only be construed as an indicator of true low risk if the patient received no targeted interventions.
+Figure 1: Stratification of inpatient admissions based on fall events and targeted interventions.
+The intervention cohort was excluded from primary optimization analyses but retained for post-optimization validation.
 We partition patient encounters into three cohorts with distinct labeling strategies:
-  • Fall cohort: All encounters with at least one fall event, regardless of targeted interventions applied. These encounters are labeled Safe High-Risk.
-  • Reference cohort: Encounters with no targeted interventions throughout the encounter, and no fall events. These encounters are labeled Safe Low-Risk.
+  • Fall cohort: All encounters with at least one fall event, regardless of targeted interventions applied. These encounters are labeled high-risk.
+  • Reference cohort: Encounters with no targeted interventions throughout the encounter, and no fall events. These encounters are labeled low-risk.
   • Intervention cohort: Encounters with at least one targeted intervention during the encounter, and no fall events.
 These encounters are considered to have uncertain risk, and are not labeled for the optimization.
-Figure 1: Stratification of inpatient admissions based on fall events and targeted interventions.
-The intervention cohort was excluded from primary optimization analyses but retained for post-optimization validation.
-The extreme imbalance in cohort sizes–with 80.7% of patients in the ambiguous intervention cohort–underscores why standard supervised learning approaches fail in this setting.
-Simply assigning these 43,799 patients to a risk category based on their JHFRAT scores or fall outcomes would introduce massive label noise, as we would be guessing the counterfactual (what would have happened without interventions) for the majority of our training data.
-Our partial supervision approach instead acknowledges this uncertainty explicitly, training only on the 10,410 patients (19.3%) for whom we have confident labels while reserving the intervention cohort for post-optimization validation to assess how the optimized models generalize to the ambiguous majority.
+The extreme imbalance in cohort sizes underscores why standard supervised learning approaches fail in this setting.
+Simply assigning the 38,643 intervention cohort patients to a risk category based on their JHFRAT scores or fall outcomes would introduce massive label noise, as we would be guessing the counterfactual (what would have happened without interventions) for the majority of our training data.
+Our partial supervision approach instead acknowledges this uncertainty explicitly, training only on the 8,348 patients (17.8% of included encounters) for whom we have confident labels while reserving the intervention cohort for post-optimization validation to assess how the optimized models generalize to the ambiguous majority.
 4.2 MIP Models
 The JHFRAT scale is divided, per current practice, into three risk categories: Low-Risk ($k=1$) for scores less than 6, Moderate-Risk ($k=2$) for scores 6-13, and High-Risk ($k=3$) for scores greater than 13.
+We apply the corresponding true labels to the labeled training data: $k^{*}=1$ for the low-risk patients in the reference cohort and $k^{*}=3$ for the high-risk patients in the fall cohort.
 While our general framework jointly optimizes both scoring weights $\beta$ and thresholds $\tau$, the JHFRAT application presents unique deployment constraints that motivate a modified approach.
@@
 This decision reflects a common practical scenario where healthcare systems require incremental refinements to existing tools rather than complete redesigns, as changing risk thresholds would necessitate updates to intervention protocols, clinical guidelines, and quality metrics across the entire health system.
-The absence of reliable moderate-risk labels in our dataset further supports this constrained optimization approach.
-Due to the selective labeling problem, we can confidently identify only extreme cases: patients who fell despite interventions (Safe High-Risk) and those who remained fall-free without targeted interventions (Safe Low-Risk).
-Patients with moderate JHFRAT scores predominantly received interventions (78.3% in our data), making their true risk indeterminate—they might be correctly classified moderate-risk patients, or they might be misclassified high or low-risk patients.
-Rather than introducing noise by guessing labels for this ambiguous middle group, we apply singleton feasible sets only to the extreme cases: $\mathcal{S}_{i}=\{1\}$ for all $i$ in the Safe Low-Risk cohort and $\mathcal{S}_{i}=\{3\}$ for all $i$ in the Safe High-Risk cohort.
-This configuration represents a special but important case of our general framework where partial supervision is extreme—no labeled examples exist for intermediate categories.
-The fixed thresholds naturally prevent the middle category collapse problem described in Section 2.3, as the moderate-risk band maintains its width of $\tau_{2}-\tau_{1}=7$ points regardless of the optimization.
-While this approach doesn’t leverage our framework’s full capability for joint score-threshold learning, it demonstrates its flexibility in handling severely limited supervision scenarios common in healthcare, where intervention effects obscure risk for the majority of patients.
-Future work could explore threshold optimization using external validation data or prospective trials where moderate-risk labels can be more reliably determined.
+We note that this fixed-threshold approach, while more restrictive than full joint optimization, offers several practical advantages: (1) it ensures the optimized tool remains immediately deployable without systemic changes to care protocols, (2) it maintains comparability with historical JHFRAT data and quality metrics, and (3) it reduces the optimization complexity, improving solution stability and interpretability.
+The resulting problem focuses on finding optimal integer weights $\beta\in\mathbb{Z}^{17}_{+}$ that best separate the extreme risk groups within the established scoring scale, effectively treating this as a constrained parameter refinement problem rather than a full model redesign.
+The complete JHFRAT-optimizing MIP formulation with these fixed thresholds is given as follows:
+$\min_{\beta,\tau,z}\quad$ $\sum_{i=1}^{n}w_{i}\sum_{k=1}^{3}\ell(k,k_{i}^{*})\cdot z_{ik}$ (11a)
+subject to $\sum_{k=1}^{3}z_{ik}=1,\quad\forall i$ (11b)
+(1d) - (1h)
+$0\leq\beta_{j}\leq 13\quad\forall j$ (11c)
+$\beta_{i}\leq\beta_{j}\quad\forall(i,j)\in P$ (11d)
+$\frac{\sum_{i\in[n]:k_{i}^{*}=3}z_{i1}}{|\{i:k^{*}_{i}=3\}|}\leq\xi_{sFNR}$ (11e)
+$\frac{\sum_{i\in[n]:k_{i}^{*}=3}z_{i2}}{|\{i:k^{*}_{i}=3\}|}\leq\xi_{mFNR}$ (11f)
+$\frac{\sum_{i\in[n]:k_{i}^{*}=1}z_{i3}}{|\{i:k^{*}_{i}=1\}|}\leq\xi_{sFPR}$ (11g)
+$\frac{\sum_{i\in[n]:k_{i}^{*}=1}z_{i2}}{|\{i:k^{*}_{i}=1\}|}\leq\xi_{mFPR}$ (11h)
+$\beta_{j}\in\mathbb{Z},\quad z_{ik}\in\{0,1\}$ (11i)
+The objective function (11a) exactly follows from (1a) by optimizing over all $n$ training samples and 3 JHFRAT categories.
+The definition for $\ell(k,k_{i}^{*})$ follows from (2), and we set $\omega_{under}=3$ and $\omega_{over}=1$ so that the ordinal loss for under-triage is three times that of over-triage.
+The sample weights $w_{i}$ are derived from the number of samples in each class such that the total contribution to the objective function is equal between the imbalanced low and high-risk classes: $w_{i}=\frac{n}{|\{j:k_{j}^{*}=k_{i}^{*}\}|}.$
 We apply three types of governance constraints in our MIP formulations for JHFRAT.
-Firstly, we apply a uniform lower and upper bound on all model coefficients.
+Firstly, we apply a uniform lower and upper bound on all model coefficients via (11c).
 Since we only include the risk factors from JHFRAT with positive coefficients, it is reasonable to impose the clinical assumption inherent in the lower bound of 0 that all of these factors can only contribute positively to high fall risk.
 Furthermore, we limit coefficients to the highest class’s lower bound of 13 for model stability.
-Secondly, we impose integer constraints on the coefficients to maintain consistency with the format of JHFRAT and ease cognitive burden of manual assessment.
-Finally, we impose monotonicity constraints for hierarchical risk factors to mantain clinical validity.
-In particular, the risk factors in the categories of age, medications, and patient care equipment are inherently hierarchical, and we let $P$ denote the set of pairs of risk factor indices corresponding to these pairwise orderings (i.e.
-the model coefficient for one high-fall risk drug is less than or equal to the coefficient for two or more high-fall risk drugs).
-The complete JHFRAT-optimizing MIP formulation is:
-$\min_{\beta,\tau,y,z}\quad$ $\sum_{i=1}^{n}w_{i}\sum_{k=1}^{3}c(k,\mathcal{S}_{i})\cdot z_{ik}$ (19a)
-subject to $\sum_{k=1}^{3}z_{ik}=1,\quad\forall i$ (19b)
-(1d) - (1h) (19c)
-$0\leq\beta_{j}\leq 13\quad\forall j$ (19d)
-$\beta_{i}\leq\beta_{j}\quad\forall(i,j)\in P$ (19e)
-$\beta_{j}\in\mathbb{Z}\quad y_{ik},z_{ik}\in\{0,1\}$ (19f)
-We consider four variations to the MIP model above:
-  1. 1.
-    Symmetric: The ordinal loss for under-triage and over-triage is the same ($\alpha_{under}=\alpha_{over}=1$). Constraints and other hyperparameters follow exactly from Model (19).
-  2. 2.
-    Asymmetric: The ordinal loss for under-triage is three times that of over-triage ($\alpha_{under}=3,\alpha_{over}=1$). Constraints and other hyperparameters follow exactly from Model (19).
-  3. 3.
-    False-Positive (FP) constrained: The ordinal loss for under-triage is three times that of over-triage, and all other hyperparameters follow exactly from Model (19).
-Let $\bar{\beta}$ indicate the JHFRAT model coefficients and let $\bar{z}$ indicate the feasible set binary classification variables imposed deterministically by the MIP formulation with $\bar{\beta}$.
-We impose the following two constraints, in addition to those in Model (19):
-    $\sum_{i\in[n]|0\in S_{i}}z_{i1}$ $\leq\sum_{i\in[n]|0\in S_{i}}\bar{z}_{i1}$ (20a)
-    $\sum_{i\in[n]|0\in S_{i}}z_{i2}$ $\leq\sum_{i\in[n]|0\in S_{i}}\bar{z}_{i2}$ (20b)
-    These additional performance constraints ensure that the number of Safe Low-Risk labeled encounters classified as Moderate-Risk (20a) and High-Risk (20b) cannot exceed the numbers of such classifications with the JHFRAT model.
-  4. 4.
-    False-Negative (FN) & False-Positive (FP) constrained: The ordinal loss for under-triage is three times that of over-triage.
-The ordinal loss for under-triage is three times that of over-triage, and all other hyperparameters follow exactly from Model (19).
-With the same definition of $\bar{z}$ as above, we impose the following four constraints, in addition to those in Model (19):
-    $\sum_{i\in[n]|2\in S_{i}}z_{i0}$ $\leq\sum_{i\in[n]|2\in S_{i}}\bar{z}_{i0}$ (21a)
-    $\sum_{i\in[n]|2\in S_{i}}z_{i1}$ $\leq\sum_{i\in[n]|2\in S_{i}}\bar{z}_{i1}$ (21b)
-    $\sum_{i\in[n]|0\in S_{i}}z_{i1}$ $\leq 0.05\sum_{i\in[n]|0\in S_{i}}1$ (21c)
-    $\sum_{i\in[n]|0\in S_{i}}z_{i2}$ $\leq 0.3\sum_{i\in[n]|0\in S_{i}}1$ (21d)
-    The number of Safe High-Risk labeled encounters classified as Low-Risk (21a) and Moderate Risk (21b) cannot exceed the numbers of such classifications with the JHFRAT model.
-The portion of Safe Low-Risk labeled encounters classified as Moderate-Risk (21c) and High-Risk (21d) cannot exceed given thresholds (30% and 5% respectively).
-These constraints allow for a small increase from the JHFRAT benchmark of 26.5% and 0.9% for these same categorizations, as seen in Table 2.
-We note that this fixed-threshold approach, while more restrictive than full joint optimization, offers several practical advantages: (1) it ensures the optimized tool remains immediately deployable without systemic changes to care protocols, (2) it maintains comparability with historical JHFRAT data and quality metrics, and (3) it reduces the optimization complexity, improving solution stability and interpretability.
-The resulting problem focuses on finding optimal integer weights $\beta\in\mathbb{Z}^{17}_{+}$ that best separate the extreme risk groups within the established scoring scale, effectively treating this as a constrained parameter refinement problem rather than a full model redesign.
+Secondly, in (11d) we impose monotonicity constraints for hierarchical risk factors to maintain clinical validity.
+In particular, the risk factors in the categories of age, medications, and patient care equipment are inherently hierarchical, and we let $P$ denote the set of pairs of risk factor indices corresponding to these pairwise orderings (i.e., the model coefficient for one high-fall risk drug is less than or equal to the coefficient for two or more high-fall risk drugs).
+Finally, in (11i) we impose integer constraints on the coefficients to maintain consistency with the format of JHFRAT and ease cognitive burden of manual assessment.
+We impose four performance constraints using the number of false positives and false negatives from the original JHFRAT model.
+The number of High-Risk labeled encounters classified as Low-Risk (11e) and Moderate Risk (11f) cannot exceed the rates of such “severe” and “moderate” false-negative classifications with the JHFRAT model, $\xi_{sFNR}$ and $\xi_{mFPR}$.
+Conversely, the number of Low-Risk labeled encounters classified as High-Risk (11g) and Moderate Risk (11h) cannot exceed the rates of the severe and moderate false-positive classifications with the JHFRAT model, $\xi_{sFPR}$ and $\xi_{mFPR}$.
+These performance constraints ensure that improvements achieved by the MIP model do not come at the expense of worse performance for one class or the other.
 4.3 Experimental Setup
-We split the encounters from the combined fall and reference cohorts into an 80% (encounters) training and validation set, and a 20% (encounters) test set, stratified by risk label.
+We split the encounters from the combined fall and reference cohorts into an 80% training set, and a 20% test set, stratified by risk label and current JHFRAT category.
 We retain the intervention cohort separately as an exogenous dataset for concordance analysis post-optimization.
-We perform 5-fold cross-validation for training the CSO, and run a hyperparameter gridsearch over the values of $\alpha_{under}$, $\alpha_{over}\in\{0.5,1,2,4\}$ and $\lambda_{1}\in\{0.2,0.5,0.8\}$ (such that $\lambda_{1}=\lambda_{1}^{-}=\lambda_{1}^{+}$ and $\lambda_{2}=1-\lambda_{1}=\lambda_{2}^{-}=\lambda_{2}^{+}$).
-For each set of hyperparameters, we consider the average coefficients across the cross-validation folds.
-Each MIP model is trained on the entire training set, utilizing the best feasible rounded coefficients from the CSO gridsearch.
-If none of the rounded CSO coefficients are feasible for a given MIP model, the JHFRAT coefficients are used for the warm-start.The CSO model is solved with cvxpy, and all MIP variations are solved with Gurobi 12.0.3.
+We first train a series of CSO models via a hyperparameter gridsearch over the values of $\alpha^{-}$, $\alpha^{+}\in\{0.5,1,2,4\}$ and $\lambda_{1}\in\{0.2,0.5,0.8\}$ (such that $\lambda_{1}=\lambda_{1}^{-}=\lambda_{1}^{+}$ and $\lambda_{2}=1-\lambda_{1}=\lambda_{2}^{-}=\lambda_{2}^{+}$).
+For each set of hyperparameters, we perform 5-fold cross-validation over the training set and average the coefficients across the cross-validation folds.
+The set of candidate warm-start solutions is the sets of these averaged coefficients, rounded to the nearest integer.
+The MIP model is trained on the entire training set, utilizing the candidate warm-start solution that has the best MIP objective value while satisfying all MIP constraints.
+If no candidate warm-start solutions are feasible, we warm-start with the original JHFRAT coefficients.
+The CSO models are solved with MOSEK via cvxpy, and all MIP variations are solved with Gurobi 12.0.3.
 4.4 Baselines and Evaluation Metrics
 We compare the outcomes of our approach against the original JHFRAT model with expert-derived weights.
-Furthermore, we compare the CSO performance with those of the MIP models in order to evaluate the added value of the MIP portion of the optimization pipeline.
+Furthermore, we compare the CSO performance with that of the MIP models in order to evaluate the added value of the MIP portion of the optimization pipeline.
 We utilize the following metrics to compare model performance:
-  • Classification Accuracy: We define tight accuracy as the portion of encounters correctly classified into low or high-risk categories.
-We further define loose accuracy to additionally include all encounters classified as moderate-risk as correctly classified.
-  • High-Risk Precision and Recall: To account for the imbalance of the dataset, we rely on the high-risk precision and recall in addition to the accuracy metrics.
-  • Binary Classification Performance Measures: We utilize the distribution of scores for the optimized models to generate receiver operating characteristic (ROC) and precision-recall curves.
-We then utilize the area under each curve, AUROC and AUPRC respectively, as model performance metrics.
+4.4.1 Classification Accuracy
+An encounter data sample is considered correctly classified only if it is exactly classified into its correct low-risk or high-risk category.
+A data sample is considered mis-classified with “moderate error” if it is classified into the moderate-risk category, and mis-classified with “severe error” if it is mis-classified into the opposite category: a low-risk labeled sample classified into the high-risk category and vice-versa.
+We define tight accuracy as the portion of encounters correctly classified, and define loose accuracy to additionally include all instances of moderate error as correctly classified.
+4.4.2 High-Risk Precision and Recall
+We rely on the high-risk precision and recall to account for the imbalance of the dataset and the importance of correctly identifying high-risk patients in this setting.
+For these binary-classification-based metrics, moderate and low-risk classifications are grouped together.
+4.4.3 Binary Classification Performance Measures
+We utilize the distribution of scores for the optimized models to generate receiver operating characteristic (ROC) and precision-recall curves.
+We then utilize the area under each curve, AUROC and AUPRC, respectively, as model performance metrics.
 5 Results
-4.1 Optimized Risk Factor Coefficients
-The optimal solutions to the four MIP variations are featured in Table 2, in comparison with the current JHFRAT scoring coefficients.
-The optimal CSO coefficients for symmetric class loss (hyperparameters $\alpha_{under},\alpha_{over}=1,\lambda_{1}^{-},\lambda_{1}^{+},\lambda_{2}^{-},\lambda_{2}^{+}=1$) are also included as a representative example of CSO results.
+The optimal solution to the MIP defined in (11) is featured in Table 2, in comparison with the current JHFRAT scoring coefficients.
+The optimal feasible CSO warm-start coefficients for this MIP (hyperparameters $\alpha^{-}=0.5,\alpha^{+}=1,\lambda_{1}^{-}=\lambda_{1}^{+}=0.8,\lambda_{2}^{-}=\lambda_{2}^{+}=0.2$) are also included.
+Both the MIP and CSO warm-start have total coefficient sums greater than the current JHFRAT, promoting higher possible total scores.
+Cognition and elimination risk factors feature far more prominently in the MIP model than in the current JHFRAT.
+Conversely, all patient care equipment coefficients are zero in the optimal MIP model.
 Table 2: Risk factor coefficients across baseline and optimized models
-Risk Factor JHFRAT CSO Symmetric Asymmetric FP-cap FN&FP-cap
-Age 60-69 years 1 1.70 1 0 1 0
-Age 70-79 years 2 1.92 2 3 1 1
-Age $\geq$ 80 years 3 2.22 2 13 4 2
-Incontinence 2 7.33 11 9 4 12
-Urgency/frequency 2 3.75 3 6 5 3
-Altered awareness 1 7.66 12 9 12 10
-Impulsive 2 15.01 13 11 12 7
-Lack of understanding 4 5.20 13 12 3 12
-One present 1 0.77 0 3 2 2
-Two present 2 1.38 1 3 3 6
-Three or more present 3 1.56 2 13 3 8
-Fall within 6 months 5 2.22 0 6 0 0
-One HFRD 3 2.60 2 2 1 0
-Two or more HFRD 5 4.04 4 3 3 0
-Sedation procedure 7 4.04 4 3 3 0
-Requires assistance 2 6.84 11 13 5 8
-Unsteady gait 2 6.76 13 10 0 7
-Visual/auditory impairment 2 6.92 6 2 9 11
-Total Coefficient Sum 49 81.92 100 121 71 89
-All optimized models have total coefficient sums greater than the current JHFRAT, and thus we compare feature contributions via their percent contributions to the total coefficient sum.
-A few trends persist across the model variations.
-Cognition risk factors feature far more prominently in the MIP models than in the current JHFRAT.
-Conversely, the history of falls risk factor only persists as non-zero in the asymmetric MIP model, albeit with less proportional importance (7.4%) compared to JHFRAT (10.2%).
-4.2 Score Distributions and Categorization
-Figure 2 reveals how different optimization objectives reshape the risk score distributions compared to the original JHFRAT.
+Category Risk Factor JHFRAT CSO MIP
+Age 60-69 years 1 1 0
+70-79 years 2 1 2
+$\geq$ 80 years 3 2 2
+Elimination Incontinence 2 6 6
+Urgency/frequency 2 3 7
+Cognition Altered awareness 1 6 13
+Impulsive 2 12 12
+Lack of understanding 4 3 4
+Equipment 1 present 1 1 0
+2 present 2 1 0
+3+ present 3 1 0
+Fall Hist. Fall within 6 months 5 2 1
+Medication 1 high fall risk drug 3 3 2
+2+ high fall risk drugs 5 4 2
+Sedation procedure 7 4 2
+Mobility Requires assistance 2 4 7
+Unsteady gait 2 4 2
+Visual/auditory impairment 2 6 3
+Total Coeff. Sum 49 64 65
+Table 3 details the model categorization accuracy per labeled class for each model.
+Both optimized models improve correct classification on the training data for both low-risk and high-risk patients while reducing or maintaining the rates of moderate and severe error, as designed.
+For the training data, correct low-risk classification increases from 71% (original JHFRAT) to 83% (MIP), and correct high-risk classification increases from 30% to 46%.
+This improvement persists in the testing data, with an increases to 84% and 48% correct classification for low-risk and high-risk patients respectively, with the JHFRAT testing data rates remaining the same.
+Furthermore, the severe and moderate error rates for the models remain the same or lower than the JHFRAT rates for the testing data.
+The severe error rates remain comparable across models, likely because the severe error performance constraints (5e) and (5g) are in direct conflict with each other: there is a natural limit, to how much the model can reduce the severe false positives without increasing severe false negatives, and vice versa.
+For the training data, these constraints are explicitly enforced, but the effects of these constraints persist in the testing data, as evident by the consistency in severe error rates for testing data.
+The optimized models outperform JHFRAT in area under the precision-recall curve for the testing datasets, per Table 4 and Figure 2.
+The MIP’s focus on categorizing around thresholds is reflected in its improved accuracy metrics over the CSO model, despite having similar AUC scores.
+Table 3: Train and test data categorization comparison across original and optimized models
+Model Label Categorization
+Correct Mod. Err. Severe Err.
+Training Data
+Orig. JHFRAT Low-Risk 4,519 (71%) 1,788 (28%) 58 (1%)
+High-Risk 105 (30%) 202 (58%) $\bm{42(12\%)}$
+CSO Low-Risk 4,940 (78%) 1,377 (22%) $\bm{48(1\%)}$
+High-Risk 151 (43%) 156 (45%) $\bm{42(12\%)}$
+Opt. MIP Low-Risk $\bm{5,271(83\%)}$ $\bm{1,036(16\%)}$ 58 (1%)
+High-Risk $\bm{160(46\%)}$ $\bm{147(42\%)}$ $\bm{42(12\%)}$
+Testing Data
+Orig. JHFRAT Low-Risk 1,135 (71%) 441 (28%) 15 (1%)
+High-Risk 26 (30%) 51 (58%) $\bm{11(13\%)}$
+CSO Low-Risk 1,279 (80%) 301 (19%) $\bm{11(1\%)}$
+High-Risk 39 (45%) 40 (45%) $\bm{9(10\%)}$
+Opt. MIP Low-Risk $\bm{1,342(84\%)}$ $\bm{235(15\%)}$ 14 (1%)
+High-Risk $\bm{42(48\%)}$ $\bm{35(40\%)}$ $\bm{11(13\%)}$
+Table 4: Performance metrics on test dataset
+Model Tight Acc. Loose Acc. ROC PRC Prec. Rec.
+Orig. JHFRAT 0.69 0.98 0.88 0.51 0.63 0.29
+CSO 0.79 $\bm{0.99}$ 0.93 0.66 $\bm{0.78}$ 0.44
+Opt. MIP $\bm{0.82}$ 0.98 $\bm{0.94}$ $\bm{0.67}$ 0.75 $\bm{0.48}$
+Figure 2: Receiver Operating Characteristic (ROC) and Precision-Recall performance curves for JHFRAT and optimized models on testing dataset.
+Both the CSO warm-start and optimized MIP improve over the original JHFRAT.
+Figure 3 reveals how the optimized model coefficients reshape the risk score distributions compared to the original JHFRAT.
 The Kolmogorov-Smirnov (KS) statistic quantifies the maximum vertical separation between the cumulative distribution functions of the Safe Low-Risk and Safe High-Risk cohorts, with larger values indicating better discriminative ability.
-The associated p-values test the null hypothesis that the two risk groups come from the same distribution—highly significant p-values (all p < 0.001) confirm that all models successfully separate the risk groups, though with varying degrees of separation.
-Figure 2: Score distributions for all labeled data (test and train data together)
-The distribution patterns directly reflect each model’s clinical optimization strategy:
-Original JHFRAT–Conservative but Imprecise: The overlapping distributions (KS = 0.620) show why current practice struggles: many true high-risk patients score in the low-to-moderate range (under-triage), while some low-risk patients score high (over-triage).
+The associated p-value tests the null hypothesis that the two risk groups come from the same distribution.
+The original JHFRAT, CSO warm-start and optimized MIP models all have highly significant p-values (all p ¡ 0.001), confirming that all models successfully separate the risk groups, though with varying degrees of separation.
+The distribution of the unlabeled intervention cohort is included for reference.
+Figure 3: Score distributions for all study data (test and train data together).
+Both CSO and MIP improve separation of high-risk and low-risk samples, at the expense of skewing high-risk and unlabeled samples to higher scores.
+The overlapping distributions in the original JHFRAT model (KS = 0.620) show why current practice struggles: many true high-risk patients score in the low-to-moderate range (under-triage), while some low-risk patients score high (over-triage).
 This overlap drives both missed falls and unnecessary interventions.
-With 44.8% of eventual fall patients classified as low or moderate risk, critical prevention opportunities are missed.
-Symmetric MIP–Maximum Separation for Clear Decision-Making: The stark bimodal distribution (KS = 0.699) eliminates the ambiguous middle ground, pushing patients toward clear low-risk or high-risk classifications.
-This suits healthcare systems that prefer definitive decisions: either minimal monitoring (score < 6) or full prevention protocols (score > 13).
-The trade-off is extreme scores (up to 60) that may seem implausible to clinicians, potentially undermining trust.
-Best for: facilities with binary intervention protocols where middle-ground treatments offer little value.
-Asymmetric MIP–Aggressive Fall Prevention: The rightward shift of both distributions reflects the 3:1 penalty for under-triage, effectively lowering the bar for high-risk classification.
-This achieves 86% sensitivity for falls (vs.
-45% in original JHFRAT) but classifies 17% of safe low-risk patients as high-risk.
-The extreme score range (up to 70) results from the optimizer pushing borderline cases strongly rightward to avoid missing falls.
-Best for: facilities where fall consequences are severe (e.g., surgical or elderly units) and resources allow broader intervention deployment.
-FP-cap MIP–Targeted High-Risk Identification: By constraining false positives to current levels, this model (KS = 0.671) maintains clinician workload while better identifying true high-risk patients.
-The distribution shows selective rightward movement—pushing true fall patients across the high-risk threshold while keeping most low-risk patients in place.
-Scores remain clinically plausible (maximum 35).
-This achieves 45% fall sensitivity with 81% positive predictive value, meaning when it flags high-risk, it’s usually correct.
-Best for: resource-constrained settings where intervention capacity is fixed and false alarms cause alert fatigue.
-FP&FN-cap MIP—Balanced Improvement Within Constraints: The bidirectional constraints produce the most JHFRAT-like distribution while still achieving better separation (KS = 0.683 vs.
-0.620).
-Both cohorts shift toward the extremes but within bounds that maintain operational compatibility.
-This model essentially "tune” JHFRAT rather than replacing it, achieving 64% fall sensitivity while keeping false positive rates manageable.
-Best for: conservative healthcare systems requiring evidence that new models won’t disrupt operations before accepting larger changes.
-CS–Smooth Risk Stratification: The continuous optimization produces graduated separation without the stark bimodality of MIP solutions.
-The smooth distribution (KS = 0.700) suggests CSO finds a natural risk continuum rather than forcing binary separation.
-This may better reflect clinical reality where risk truly is continuous, not discrete.
-With 66% fall sensitivity and moderate false positives, it balances competing objectives.
-Best for: initial implementations where extreme changes might face resistance, or as a warm-start for subsequent MIP refinement.
-The key insight is that no single distribution is universally "best"–each reflects different clinical priorities and operational constraints.
-The original JHFRAT’s overlapping distributions explain its poor performance, while our optimized models offer a menu of alternatives: maximum separation (Symmetric MIP) for clear decisions, protective bias (Asymmetric MIP) for high-stakes environments, resource-aware precision (FP-cap) for efficiency, or incremental improvement (FP&FN-cap) for conservative adoption.
-The extreme scores in unconstrained models, while mathematically optimal, highlight the importance of including clinical face-validity constraints in future deployments.
-We calculate the score differentials (MIP score - JHFRAT score) for all encounters in the dataset, as seen in Figure 3.
-All MIP models results in a positive average score differential, with a heavy upper tail.
-The asymmetric MIP without any performance constraints results in the most positive skew and the FP-cap MIP results in the least positive skew.
-Figure 3: Score differentials (MIP score - JHFRAT score) for all patients in dataset
-Skewness in the score differentials is also reflected in the categorization of encounters across the low, moderate and high-risk categories shown in Table 3.
-Without additional performance constraints, both the symmetric and asymmetric models significantly increase the number of encounters categorized as high-risk in the test data set.
-The symmetric model results in a 1100% increase in high-risk false positives and a 164% increase in true positives while the asymmetric model has a 1878% increase in high-risk false positives and 200% increase in true positives in the test set.
-The two models with false positive maximums still achieve notable gains in true positives: 56% and 124% increases from JHFRAT for FP-cap and FN&FP-cap, respectively.
-Both models also decrease the number of safe high-risk patients categorized as moderate-risk, but increase the number categorized as low-risk.
-All models except for the asymmetric MIP without performance constraints achieve a small increase in the number of true negative safe low-risk patients identified as low-risk.
-Table 3: Test data encounter categorization across baseline and optimized models
-Model Encounters Model Categorization
-Low-Risk Moderate-Risk High-Risk
-Original JHFRAT Safe Low-Risk 1,149 (72.6%) 420 (26.5%) 14 (0.9%)
-Safe High-Risk 7 (8.1%) 55 (63.2%) 25 (28.7%)
-All 17,007 (31.4%) 30,254 (55.8%) 6,946 (12.8%)
-CSO Safe Low-Risk 1,156 (73.0%) 374 (20.5%) 53 (3.3%)
-Safe High-Risk 8 (9.2%) 22 (25.3%) 57 (65.5%)
-All 15,640 (28.9%) 18,124 (33.43%) 20,443 (37.7%)
-Symmetric MIP Safe Low-Risk 1,180 (74.5%) 262 (16.6%) 168 (10.6%)
-Safe High-Risk 7 (8.1%) 14 (16.1%) 66 (75.9%)
-All 16,207 (29.9%) 10,581 (19.5%) 27,419 (50.58%)
-Asymmetric MIP Safe Low-Risk 946 (59.8%) 360 (22.7%) 277 (17.5%)
-Safe High-Risk 6 (6.9%) 6 (6.9%) 75 (86.2%)
-All 11,278 (20.8%) 9,212 (17.0%) 33,717 (62.2%)
-FP-cap MIP Safe Low-Risk 1,258 (79.5%) 316 (20.0%) 9 (0.6%)
-Safe High-Risk 12 (13.8%) 36 (41.4%) 39 (44.8%)
-All 18,142 (33.5%) 22,892 (42.2%) 13,172 (24.30%)
-FP&FN-cap MIP Safe Low-Risk 1,237 (78.1%) 294 (18.6%) 52 (3.3%)
-Safe High-Risk 11 (12.6%) 20 (23.0%) 56 (64.4%)
-All 18,034 (33.3%) 15,562 (28.7%) 20,611 (38.0%)
-4.3 Performance Metrics
-Table 4 includes the performance metrics calculated for each model.
-In general, the optimized models achieve moderate gains in tight accuracy at the expense of loose accuracy.
-Only the FP-cap MIP improves on the tight accuracy without loss of loose accuracy.
-Table 4: Performance metrics on test dataset
-Model Accuracy AUC High-Risk Precision High-Risk Recall
-Tight Loose AUROC AUPRC
-Original JHFRAT 0.70 0.99 0.92 0.51 0.64 0.29
-CSO 0.73 0.96 0.94 0.65 0.52 0.66
-Symmetric MIP 0.75 0.90 0.94 0.65 0.28 0.76
-Asymmetric MIP 0.61 0.83 0.92 0.57 0.21 0.86
-FP-cap MIP 0.78 0.99 0.92 0.62 0.81 0.45
-FN&FP-cap MIP 0.77 0.96 0.90 0.63 0.52 0.64
-All optimization models outperform JHFRAT in area under the precision-recall curve for the testing and training datasets.
-All models perform similarly with high recall in the testing dataset, with the curves only notably separating at recall below 0.6.
-In contrast, the optimized models achieve visibly higher precision across all recall points for the training data.
-The small improvements in AUROC achieved by the models in the training set do not materialize in the testing set.
-Figure 4: Receiver Operating Characteristic (ROC) and Precision-Recall performance curves across JHFRAT and five optimization models on training dataset (above) and testing dataset (below).
+Both the CSO warm-start and optimized MIP improve on the KS statistic, at the expense of skewing the high-risk and unlabeled distributions rightward and increasing the maximum total scores from 33 for the original JHFRAT to 50 and 55 for CSO and MIP, respectively.
 6 Discussion
-5.1 Key Findings
-The proposed CSO and mixed-integer programming pipeline achives significant performance gains compared to the current JHFRAT model while offering priority-based tuning options in the hyperparameters.
-In general, the score distributions and sample categorizations respond as expected to the provided hyperparameters.
-Utilizing a symmetric distance function for mis-classification loss results in nearly symmetrical mis-classification rates for low-risk and high-risk labeled samples, while the use of an asymmetric 3-to-1 high-risk to low-risk distance function results in higher skew of both classes towards high-risk categorization.
-The notable exception to expectations is the increase in low-risk false negatives among the two performance-constrained models.
-The false-negative constraint for the FN&FP-cap MIP should mitigate an increase in such classifications, but this cannot be guaranteed in the testing setting.
-While the CSO and MIP models themselves do not inherently discourage moderate-risk classification, when applied in a setting with no such moderate risk labels, the resulting score distributions disaggregate from the mean.
-This movement away from moderate risk classification can come at the expense of increasing further-class mis-classification, as is the case for the FP-cap and FN&FP-cap models.
-Nevertheless, these models remain promising options for improving true high-risk classification without drastic increases in the total number of patients classified as high-risk.
-5.2 Limitations and Future Work
-Several model extensions warrant investigation:
-  • Non-linear scoring functions: Extending to tree-based or shallow neural architectures while maintaining interpretability
-  • Time-varying risk: Incorporating temporal dynamics in sequential assessment settings
-  • Multi-site learning: Federated optimization across institutions with heterogeneous populations
-There are a few key limitations with the JHFRAT model experiments.
-Use of average JHFRAT scores over the encounter for model training results in some non-binary risk score values.
-These binary risk-score values represent averages over the encounter, but may contribute to extreme values in the optimized coefficients (i.e.
-coefficients larger than the highest class lower bound) in the absence of enforced upper bounds.
-The exclusion of all intervention cohort encounters also potentially limits the generalizability of the optimized models.
-Future work could explore including some patients from the intervention cohort as Safe Moderate-Risk, or apply multiple labels, depending on the number and type of fall prevention interventions applied.
-5.3 Clinical Deployment Considerations
-It is of particular clinical interest to improve identification of patients at high risk for falls while avoiding an overall increase in risk scores.
-From a resource perspective, it is impractical and wasteful to provide fall prevention interventions to patients who do not need them.
-Furthermore, from a clinical perspective, restricting mobility is an effective fall prevention method, but often leads to detrimental functional decline.
-Therefore, placing such restrictions on low-risk patients is not only unnecessary, but is actively harmful.
-These types of concerns over false positives extend beyond fall risk assessment.
-Alert fatigue is a persistent problem that is exacerbated by many false positives in assessment of risks of adverse events such as sepsis.
-The inclusion of performance constraints and the user-tunable distance function for mis-classification loss serve as tools to control the balance between under-classification and over-classification.
-Specific constraints on the number of false positives and/or false negatives for the most extreme classes offer assuredness.
-In general, the specific features of the proposed MIP formulation that are applied to any given setting can be based on the context of the risk factors, adverse event in question, available data, and current clinical practice.
-For example, enforcing optimized coefficients to be integers may be unnecessary if assessments are completed automatically and clinicians only observe the final score.
-Allowing negativity in model coefficients allows for the inclusion of risk-preventative factors in the model.
+The proposed framework makes several methodological contributions to the optimization of clinical risk scoring tools.
+First, the formulation addresses a common challenge in clinical settings where reliable labels exist only for extreme risk categories.
+When optimizing with only these extreme labels, the model may collapse intermediate thresholds to eliminate unlabeled middle categories, effectively producing a binary classifier rather than the desired ordinal tool.
+The framework prevents this degeneracy through minimum threshold gap constraints, maintaining clinically meaningful risk stratification without requiring labeled examples for every category.
+Second, the asymmetric distance-aware loss function directly encodes clinical priorities into the optimization objective.
+By allowing differential penalties for under-triage versus over-triage that scale with ordinal distance, the framework aligns the training signal with deployment consequences, which is preferable to learning symmetric costs and adjusting thresholds post-hoc.
+Third, the governance and performance constraints provide mechanisms for encoding domain knowledge and ensuring operational safety, from sign restrictions that maintain clinical face validity to false-positive and false-negative caps that guarantee minimum performance levels.
+Finally, the two-phase optimization pipeline combines the computational efficiency of convex optimization with the exactness of mixed-integer programming, enabling tractable solutions for moderate-sized clinical datasets.
+The JHFRAT case study validates these methodological contributions empirically.
+Training only on the 8,348 patients with confident extreme labels while excluding the 38,643 with intervention-censored outcomes yields models that generalize well to held-out test data, with tight accuracy improving from 69% to 82% and high-risk recall nearly doubling from 29% to 48%.
+The asymmetric loss function (3:1 under-triage to over-triage penalty) successfully biases the optimization toward high-risk identification without sacrificing overall discrimination, as evidenced by the AUPRC improvement from 0.51 to 0.67.
+The two-phase pipeline demonstrates practical value: the CSO warm-start achieves 79% tight accuracy on its own and provides feasible integer solutions that accelerate MIP convergence, while the MIP refinement yields additional gains in threshold-based classification.
+The similar AUC scores but improved accuracy metrics between CSO and MIP suggest that while both methods learn comparable score rankings, the MIP’s discrete optimization produces coefficients better calibrated to the fixed threshold boundaries.
+Beyond performance improvements, the optimized coefficients reveal clinically meaningful patterns that illustrate the value of data-driven calibration.
+Cognition-related factors, particularly altered awareness and impulsivity, receive substantially higher weights in the optimized model, increasing from 1–2 points to 12–13 points.
+Elimination-related factors similarly receive increased weights, aligning with research on bathroom-related ambulation as a high-risk period.
+Conversely, patient care equipment coefficients are reduced to zero, possibly reflecting confounding in the original design where equipment presence correlates with closer monitoring or serves as a proxy for illness severity captured elsewhere.
+These discoveries would not emerge from simply re-thresholding the original scores; they require the coefficient optimization that the framework enables.
+Several limitations warrant acknowledgment.
+The use of encounter-averaged JHFRAT scores introduces non-integer feature values, which may contribute to the large coefficient magnitudes observed; future work could optimize directly on assessment-level data or impose tighter coefficient bounds.
+The complete exclusion of the intervention cohort, while methodologically principled, may limit generalizability to patients who would typically receive interventions.
+Incorporating partial labels based on the specific interventions applied or otherwise implementing multi-label formulations could address this.
+The fixed-threshold constraint, while practically motivated for deployment compatibility, prevents full exploitation of the joint optimization framework.
+Additionally, the widened score range (maximum scores increasing from 33 to 56) may affect clinician interpretation of within-category gradation, requiring communication and training for deployment.
+The framework applies broadly to clinical risk stratification settings sharing the key structural features: ordinal categories from point-based scores, labels available only at extreme categories, and asymmetric misclassification costs.
+Examples include pressure injury risk (Braden Scale), venous thromboembolism risk (Caprini Score), and early warning scores for deterioration.
+Future extensions include non-linear scoring functions maintaining interpretability, temporal dynamics for time-varying risk, and federated optimization across institutions with heterogeneous populations.
 7 Conclusion
-In this work we develop a multi-class mixed integer programming ordinal classification model and a smooth, convex constrained score optimization.
-These models can apply in a variety of risk assessment contexts, and are particularly well suited for settings with uncertain ground truth or predictive performance requirements compared to an existing model.
-We utilize the proposed MIP and CSO formulations to conduct data-driven score optimization of the Johns Hopkins Fall Risk Assessment Tool.
-Across different model variations, we are able to increase identification of patients who fell as high-risk, and control for false positives via model parameters and constraints.
-Future work can explore refining or adapting the proposed models to other risk assessment settings.
+This work presents a mixed-integer programming framework for optimizing clinical risk scoring tools in settings where reliable labels exist only for extreme risk categories.
+The framework addresses three challenges that impede standard supervised learning in clinical deployment: missing labels for intermediate categories that can cause threshold collapse, asymmetric misclassification costs that vary with ordinal distance, and interpretability requirements that mandate integer coefficients amenable to manual calculation.
+Minimum threshold gap constraints maintain clinically meaningful risk stratification without requiring labeled examples for every category, while the constrained score optimization relaxation provides an efficient warm-start strategy that accelerates MIP convergence.
+Applied to the Johns Hopkins Fall Risk Assessment Tool, the optimized model improves high-risk recall from 29% to 48% and area under the precision-recall curve from 0.51 to 0.67, while maintaining the existing three-tier threshold structure for seamless integration with current clinical workflows.
+The optimized coefficients reveal that cognition and elimination factors are substantially more predictive than their original weights suggest, while patient care equipment contributes minimal independent information.
+These insights may inform future tool revisions beyond the current optimization.
+The framework’s flexibility in accommodating governance constraints, performance guarantees, and minimal-edit requirements positions it for broader application across clinical risk assessment domains where data-driven calibration of legacy tools is needed.
+Future extensions may incorporate time-varying risk dynamics, multi-site federated learning, and richer label structures that leverage intervention cohort information more directly.